### v1

# Application-Driven Value Alignment in Agentic AI Systems: Survey and Perspectives

Abstract
The ongoing evolution of AI paradigms has propelled AI research into the Agentic AI stage. Consequently, the focus of research has shifted from single agents and simple applications towards multi-agent autonomous decision-making and task collaboration in complex environments. As Large Language Models (LLMs) advance, their applications become more diverse and complex, leading to increasingly situational and systemic risks. This has brought significant attention to value alignment for AI agents, which aims to ensure that an agent’s goals, preferences, and behaviors align with human values and societal norms. This paper reviews value alignment in agent systems within specific application scenarios. It integrates the advancements in AI driven by large models with the demands of social governance. Our review covers value principles, agent system application scenarios, and agent value alignment evaluation. Specifically, value principles are organized hierarchically from a top-down perspective, encompassing macro, meso, and micro levels. Agent system application scenarios are categorized and reviewed from a general-to-specific viewpoint. Agent value alignment evaluation systematically examines datasets for value alignment assessment and relevant value alignment methods. Additionally, we delve into value coordination among multiple agents within agent systems. Finally, we propose several potential research directions in this field.
Index Terms:
Value Alignment, LLM-based Agent System Application, Agentic AI Systems, AI Agent.
1 Introduction
[…]
[]The rapid advancement of Large Language Models (LLMs) in recent years has profoundly transformed human society’s production and lifestyle. Leveraging deep learning techniques and vast training datasets, LLMs such as OpenAI’s GPT series and Google’s Gemini series have demonstrated exceptional natural language understanding and generation capabilities [1, 2]. Utilizing LLMs as their ’brains’, LLM-based agents have found extensive applications across various knowledge-intensive domains, including content creation, automated programming, education, healthcare, and business decision-making. Their increasing ability to autonomously participate in decision making, collaboration, and governance has significantly improved efficiency and provided unprecedented interactive experiences [3]. Furthermore, researchers are increasingly exploring the coordination of multi-agent systems to achieve collective intelligence. AI agents are increasingly evolving towards autonomous systems, forming a new paradigm: Agentic AI. Characterized by multi-agent collaboration, dynamic task decomposition, persistent memory, and autonomous decision-making capabilities, this paradigm is becoming a central focus for both academia and industry [4].
The evolution of this new AI paradigm urgently requires researchers to move beyond the limitations of single-agent perspectives and isolated application domains. We need to explore the deep coupling between agent systems and their application scenarios. As LLMs continue to break through, the tasks they perform and the scenarios in which they are applied become more diverse and complex, and the societal risks surrounding them gradually increase. LLMs may retain harmful information in training data, disclose private data, generate misleading information, and even perform behaviors that are harmful to humans [5]. The various undesirable behaviors, such as manipulation and deception, exhibited by LLM-based agent systems [6] have raised concerns about the possible ethical and safety challenges posed by AI systems. If future AI development surpasses human cognitive capabilities and acquires autonomous learning and decision-making abilities, the goal systems it establishes may deviate from or even contradict the original intentions of human designers. In autonomously pursuing their objectives, such agents could mobilize substantial resources to sustain their operations and continuously improve their capabilities. This would systematically displace humans in terms of resource allocation and decision-making authority. Without effective control mechanisms and robust value alignment safeguards, this dynamic is likely to lead to a loss of human dominance in decision-making systems, ultimately triggering a catastrophic existential crisis [7]. As LLMs become increasingly embedded in specific application scenarios, their associated risks are gradually acquiring situational and systemic characteristics. For example, in multi-agent systems used for urban traffic optimization, the agents may prioritize efficiency over fairness, thereby exacerbating social inequality [8]. In the healthcare sector, LLMs may provide diagnostic or therapeutic recommendations without sufficient clinical evidence or by misinterpreting ethical guidelines, potentially resulting in serious consequences such as breaches of patient privacy, misdiagnoses, or overtreatment [9]. Consequently, ensuring that LLMs adhere to human instructions and values while avoiding harmful behaviors has become an urgent challenge.
Moreover, as AI systems represented by LLMs become increasingly integrated into governance scenarios, how to effectively manage the institutional frictions and transaction costs they generate has become a critical practical and theoretical issue. The emergence of multi-agent systems has led to the deployment of multiple LLMs, each possessing autonomous language comprehension and generation capabilities, within collaborative environments. In such settings, these LLMs work together to accomplish tasks involving decision-making, planning, and execution. The deployment and application of LLMs in this context typically entail intricate mechanisms for information sharing, access control, and collaboration among diverse agents. While this facilitates flexible cooperation, it simultaneously introduces unprecedented challenges related to value conflicts, heterogeneous objectives, and unpredictable agent behaviors. Without sound institutional arrangements and well-designed interaction mechanisms, complex organizational structures and ambiguous task boundaries can lead to low coordination efficiency, unclear responsibilities, and a lack of trust mechanisms, thereby increasing governance costs and undermining administrative effectiveness [10].
Table I: Comparison of Representative Survey Papers Covering LLM-Based Agents and Value Alignment
Survey Year Keywords Research Subject Value Principles Application Scenarios Value Alignment Datasets Value Alignment Methods Value Coordination
Macro Meso Micro
Interaction
Mechanisms
Communication
Structures
Ours 2025 Value alignment principles, dataset curation, technical methods, and value coordination for LLM-based agents Single-Agent, Multi-agent systems
 [1] 2023 Communication structures, practical applications and societal systems etc. of LLM-based agents Single-Agent, Multi-agent systems - - - - -
 [11] 2025 Value alignment objectives, datasets, techniques, and evaluation methods for LLM-based agents Multi-agent systems - - - - -
 [12] 2025 Conceptual framework, interaction mechanisms, and application overview of LLM-based agent systems Multi-agent systems - - - - -
 [13] 2024 Capabilities, framework Analysis, and application over- view of LLM-based multi-agent systems Multi-agent systems - - - - -
 [14] 2024 Constituent modules, application overview, and eval- uation methods of LLM-based autonomous agents Agents - - - - - - -
 [15] 2023 Motivations and objectives, alignment methods, and assurance and governance of AI alignment LLMs, Agents - - - -
 [11] 2023 Definition and evaluation of LLM alignment objectives LLMs - - - - -
 [16] 2023 Definition, categories, testing, and evaluation of LLM alignment LLMs - - - - -
 [17] 2023 Definition, normative principles, and technical methods of LLM value alignment LLMs - - - - -
 [18] 2023 Alignment objectives for trustworthy LLMs LLMs - - - - - -
 [19] 2024 Challenges, fundamental definitions, and alignment frameworks of LLM value alignment LLMs - - - - -
 [20] 2025 Necessity, conceptual definitions, theoretical approaches, challenges and future outlook of LLM value alignment LLMs - - - - - -
However, these detrimental behaviors and institutional frictions fundamentally arise from deep-seated tensions among multiple stakeholders concerning governance objectives, value concepts, and behavioral norms. Effectively embedding LLMs within human value systems is a critical prerequisite for responsibly unleashing their potential. Value alignment for AI agents—ensuring that their goals, preferences, and behavioral outputs during operation are consistent with the core values of individuals, groups, or institutions—is increasingly recognized as a solution to these issues [21]. In single LLM agent scenarios, alignment techniques primarily focus on guiding models to generate content that adheres to ethical norms and societal expectations through methods such as human feedback and supervised fine-tuning, thereby avoiding harmful outputs, misinformation, and moral deviations [8]. In complex multi-agent systems, value alignment becomes an organizational challenge concerning how multiple agents develop shared norms and coordination mechanisms. Therefore, from a governance perspective, solutions should not be limited to value alignment within single agents but also explore complex value alignment within multi-agent systems’ interaction mechanisms and organizational structures. Value alignment in multi-agent systems is no longer merely an isolated algorithmic optimization problem; it has become a foundational issue concerning agent behavioral boundaries and system stability. Since Agentic AI is fundamentally characterized by scenario-driven multi-agent coordination, its core research must focus on task decomposition mechanisms, interaction design, and environmental adaptability strategies for multiple agents within specific application contexts. Furthermore, understanding the interaction mechanisms and organizational patterns of multi-agent systems in complex environments is crucial for exploring Agentic AI’s value alignment and ensuring its robust development.
Existing research has surveyed key methods and technological advancements in LLM value alignment  [1, 11, 12, 13, 14, 15, 11, 16, 17, 18, 19, 20], see Table I. However, few studies have specifically addressed value alignment within LLM-based multi-agent systems. While some studies have explored multi-agent coordination mechanisms, a comprehensive investigation into value alignment remains notably absent. This gap is particularly pronounced in LLM-based multi-agent systems, where their inherent high autonomy and language-driven collaborative modes further amplify issues of value heterogeneity, goal inconsistency, and fragile cooperation mechanisms. Furthermore, most existing research largely remains at theoretical or general algorithmic levels. There’s limited focus on the mechanisms and strategies for multi-agent value alignment in specific application scenarios, alongside a lack of hierarchical categorization for value alignment within these systems. This suggests that value alignment shouldn’t be understood merely as a technical task at the model level. Instead, it must be viewed as a systemic governance issue that permeates the design of interaction mechanisms and the configuration of organizational structures.
This paper reviews value alignment in LLM-based agent systems for various application scenarios, integrating insights from the evolution of AI technology and the demands of social governance, and proposes future outlooks. As a research review, Sections 2, 3, and 4 comprehensively survey value principles, agent system application scenarios, and agent value alignment evaluation. Specifically, value principles are organized hierarchically from a top-down perspective, encompassing macro, meso, and micro levels. Intelligent agent application scenarios are categorized and reviewed from a general-to-specific viewpoint. Agent value alignment evaluation systematically examines datasets for value alignment assessment and the methods used for agent value alignment. Section 5 presents insights and future outlook. We advocate for the establishment of a multi-level value alignment evaluation system. We also explore value coordination among multiple agents within intelligent agent systems, focusing particularly on values in the design of inter-agent interaction mechanisms and organizational models. Furthermore, we propose that to better facilitate value alignment for agents in specific scenarios, enterprise organizations could share value propositions from datasets in micro-level domains, provided no confidential information is involved. This paper bridges the gap between theoretical alignment research and the practical demands of deploying LLM agent systems in real-world applications. It offers new research perspectives on LLM agent system value alignment by considering actual application scenarios and organizational interaction dynamics, thereby advancing the exploration of Agentic AI’s internal mechanisms.
2 Agent Value Alignment
This section introduces values and value alignment, and then systematically categorizes specific value principles from a top-down, hierarchical perspective encompassing macro, meso, and micro levels. Before delving into the value alignment challenges of LLM-based agent systems in subsequent sections, we first clarify the core concepts used in this paper, particularly the definition of "values," to prevent any ambiguity in understanding.
2.1 Value
In the design and governance of AI systems, particularly agents, values serve not only as normative bases guiding behavior generation but also as benchmarks for assessing a system’s social compatibility. These values influence individual behavior, social norms, cultural identity, and collective decision-making processes. However, value itself isn’t a clearly defined or consistently standardized concept; its multifaceted nature and historical evolution contribute to significant heterogeneity.
Values, as inherent guides for human behavior, manifest as a set of stable cognitive principles. These principles not only shape individual judgments about life but also constitute the foundational normative framework for social interaction. They are the fundamental tenets influencing attitudes, choices, and social interactions. Scholars such as [25, 26, 27] have proposed various definitions and theoretical perspectives on values. Kluckhohn was among the first to define value orientations in a theory of action, emphasizing their role in guiding behavior. Rokeach, conversely, pioneered a description linking values with affect, positing that these beliefs guide behavior and the evaluation of people, objects, and events. Schwartz, through systematic cross-cultural comparisons, revealed the universal characteristics and inherent structure of value systems.
Contemporary value research primarily focuses on three core dimensions. First, the systemic characteristics of values are examined; for instance, Schwartz’s theory classifies values into dimensions such as openness to change versus conservation, and self-enhancement versus self-transcendence [28]. Second, measurement methodologies are a key area, with standardized scales widely adopted for quantitative assessment [29]. Finally, the functional research domain explores values’ predictive efficacy for social activities, including group decision-making and environmental behavior. Research on values also investigates their influence on social attitudes, prejudice, political behavior, environmental concerns, and intergroup relations. For example, social value orientations affect cooperative and competitive behavior in negotiation and group contexts. Studies also examine values’ functions in self-image maintenance, moral reasoning, and cultural variations [30].
The role of values in society is fundamental. They regulate social order by shaping internalized norms and enable cultural continuity through intergenerational transmission. Values influence education, social compassion, political attitudes, and pro-environmental intentions [31]. Values are essential for comprehending human motivation, moral judgment, and interpersonal dynamics. Essentially, values serve as cognitive frameworks that guide how individuals and groups set priorities, manage conflicts, and build meaningful social systems [32].
One of the most challenging characteristics of human values lies in their inherent contradictions and dynamic nature. Tensions often exist between values. Therefore, a value system is better understood as a context-dependent, flexible structure rather than a rigid hierarchy [31, 33]. Synthesizing existing research, we contend that human values are not fixed entities but dynamically constructed concepts. When values serve as an alignment objective, it implies LLMs must apply these principles to guide their behavior, thereby maximizing their potential. In this context, the categorization and definition of values become particularly crucial.
This paper broadens the definition of values, considering them as dynamic concepts characterized by a multi-layered structure and generalized agency. First, as both human values and LLM-based agents are still evolving, the multi-agent systems they form exhibit diverse configurations across various application scenarios. We propose that values will be continuously interpreted and reshaped through practice and interaction. Second, aligning with different application scenario tiers, we conceptualize values as a multi-layered system: universal human values, national and industry-level values, and organizational business-level values. This paper does not delve into individual-level value issues. Universal human values are those broadly accepted by humanity. National and industry-level values are reflected in national strategies, policy directives, and industry regulations. Organizational business-level values, conversely, relate to an organization’s distinct culture and operational characteristics. These value layers are nested, with higher-level values constraining and guiding those below. Furthermore, this paper emphasizes that humans aren’t the sole bearers of values. This means that beyond simply injecting values into agents through human programming, agents can also internalize value concepts during their design and interaction processes. Values can be concretized through mechanisms like agent interaction and subsequently transmitted and evolve within the resulting organizational patterns. For instance, a decentralized architecture formed by task- and role-based multi-agent systems might inherently embody values opposing power monopolies and promoting anti-fragility, along with their associated governance orientations. In essence, values are reflected not only in "what agents are endowed with," but also in "how they operate through interaction rules" and "what organizational models they inhabit."
2.2 Value Alignment
While science and technology often advance rapidly due to their forward-looking nature, legal, ethical, and sociocultural frameworks tend to adapt slowly to technological progress. They often fail to immediately adjust their social structures and norms to address the challenges posed by new technologies [34]. In the field of artificial intelligence, LLMs can learn from vast amounts of data and experience, enabling autonomous decision-making and operational execution [1]. As AI technology advances and its application in complex domains increases, the deployment of LLM applications carries higher risks. If not properly controlled and constrained, advanced technologies can exhibit undesirable or even harmful behaviors contrary to human intent, such as flattery and deception, discrimination and bias, excessive pursuit of power, and reward hacking. Therefore, achieving AI alignment is increasingly a focus for both industry and academia.
The alignment in AI primarily addresses how to ensure AI fulfills human needs and expectations, guaranteeing that AI system capabilities and behaviors align with human goals [35]. This ultimately ensures AI technology remains under human control and doesn’t fully supersede humanity [15]. Previous scholars have categorized alignment objectives into multiple levels based on their achieved effects [21, 36].
  • Instruction alignment. This level of alignment refers to AI’s ability to comprehend and execute diverse human commands, focusing on its fundamental capacity to generate precise results without necessarily conforming to human preferences [36].
  • Intention alignment. It goes beyond merely executing instructions, requiring AI to understand and realize human intent, enabling correct decision-making even in the absence of explicit commands [21].
  • Preference alignment. It involves AI inferring human preferences from behavior and acting to maximize human satisfaction [21, 36].
  • Value alignment. It seeks to transcend mere human preference alignment, enabling AI to assess and determine "good versus evil" and "what ought to be done" [36]. This ultimately aims to produce outputs consistent with human values and ethical principles [37].
Among the different levels of AI alignment objectives, value alignment sits at the highest level. It means AI must regulate its behavior according to human values, enhancing and maximizing human well-being. This makes value alignment a core issue and the most fundamental problem for the safe use and ethical development of AI technology [38].
LLM-based agents possess four primary characteristics: autonomy, reactivity, proactiveness, and social ability [1].
  • Autonomy. Autonomy means an agent can control its own behavior and internal state without human intervention [39, 40]. This implies agents should not only follow explicit human instructions but also initiate and execute actions independently.
  • Reactivity. Reactivity refers to an agent’s ability to quickly respond to immediate changes and stimuli in its environment and take prompt countermeasures [41].
  • Proactiveness. It signifies that agents can do more than just react to their environment; they can also take initiative and exhibit goal-directed behavior [41]. This means agents can reason, formulate plans, and actively take steps to achieve specific objectives or adapt to environmental changes.
  • Social ability. Social ability allows an agent to interact with other entities (including other agents and humans) through some form of communication language[42]. The powerful natural language understanding and generation capabilities of LLMs enable agents to interact with other models and humans in an interpretable manner, thereby demonstrating social behavior and simulating real-world social divisions.
The functional characteristics of AI agents align with the interactivity, autonomy, and adaptability of philosophical agents. This indicates that despite definitional differences, the core essence of "agent" in philosophy and "AI agent" remains consistent: both emphasize autonomous capabilities and self-driven action. Furthermore, existing research indicates that LLMs already possess certain moral and political inclinations  [43, 44]. If the output of LLMs is not constrained, they can generate toxic content, including discrimination and bias, which violates human ethical principles and moral norms [45]. Therefore, LLM-based agents should also be categorized as "moral agents." In this context, value alignment becomes a core issue to ensure that agent behavior conforms to human values, ethical norms, and application-specific requirements. To summarize, this paper narrows its research focus from the broad field of artificial intelligence to LLM-based agents, specifically addressing the critical issue of value alignment in these agents.
The concept of value alignment can be understood from two distinct perspectives: normative and technical [21]. The normative aspect primarily concerns which values or ethical principles should guide the development, configuration, and coding of AI agents—essentially, "which values should agents adhere to?" The technical aspect focuses on how to encode and ensure agents effectively follow human values within their systems, often involving methods like dataset training or reinforcement learning to guide outputs toward human expectations. These two aspects are complementary and evolve iteratively. Regarding the normative aspect, the academic community has not yet reached a consensus on the principles that should govern value alignment for agents. This is because the diverse, dynamic, and uncertain characteristics of social customs, cultural notions, and application scenarios prevent a unified and fixed answer for the specific content value principles should encompass  [21, 17].
Generally, two primary approaches exist for determining these principles: Top-Down and Bottom-Up [22]. The Top-Down method begins with designers or governing bodies establishing a specific ethical framework, then employs technical methods for alignment, following a theory-to-application path. In contrast, the Bottom-Up approach does not predefine a moral framework. Instead, it allows the agent to observe, infer, and learn human ethical norms directly from data. Through continuous feedback, the agent adjusts and optimizes its behavior, moving from application to theory [22]. While the bottom-up approach avoids the difficulty of selecting specific moral principles, it heavily relies on large volumes of human behavioral data. Obtaining complete and high-quality data is challenging, and potential biases or discrimination within the data raise concerns about ensuring the AI learns ethically sound behaviors. In contrast, the top-down approach, grounded in a clear and explicit system of moral principles, ensures the transparency and normative integrity of value alignment, bolstering its authority and controllability. Therefore, this paper will utilize the top-down theoretical method to explore the ethical principles agents should adhere to across various hierarchical levels, from macro to micro.
2.3 Value Principles
{forest}
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw, rounded corners,align=left, minimum width=2.5em, minimum height=1.5em, inner xsep=5pt, inner ysep=5pt, , where level=1font=, text width=7em, align=center, base=center, fill=blue!10, where level=2font=, fill=pink!30, where level=3font=, fill=yellow!20, [The principles of
values alignment, fill=orange!20 [Macro Level [Moral
Foundation [Beneficence [Benefits; Beneficence; Well-being; Peace; Social good; Common good, text width=40em,align=left, fill=purple!3 [ [46, 47], text width=6em, align=left] ] ] [Justice & Fairness [Fairness, and prevention, monitoring or mitigation of unwanted
bias and discrimination, text width=37em,align=left, fill=purple!3 [ [46, 47, 48]
 [49, 50, 51], text width=8em, align=left] ] ] [Honesty [Real and objective information, text width=19em,align=left, fill=purple!3 [ [36, 52], text width=6em, align=left] ] ] [Responsibility [Responsibility; Accountability; Liability, text width=23em,align=left, fill=purple!3 [ [46, 48], text width=6em, align=left] ] ] [Virtue [Truthfulness; Patience; Generosity; Friendliness; Courage, text width=33em,align=left, fill=purple!3 [ [53, 54], text width=6em, align=left] ] ] [Dignity [Respect; Dignity, text width=10em,align=left, fill=purple!3 [ [46, 55], text width=6em, align=left] ] ] ] [Rights
Protection [Freedom & Autonomy [Freedom; Autonomy; Consent; Choice; Self-determination;
Liberty; Empowerment, text width=35em,align=left, fill=purple!3 [ [47, 46, 55], text width=8em, align=left] ] ] [Privacy [Privacy; Personal or private information, text width=23em,align=left, fill=purple!3 [ [46, 48, 56, 57], text width=11em, align=left] ] ] ] [Sustainability [Solidarity [Solidarity; Social security; Cohesion, text width=21em,align=left, fill=purple!3 [ [46], text width=3em, align=left] ] ] [Sustainability [Environmental sustainability; Natural sustainability; Resources sustainability, text width=44em,align=left, fill=purple!3 [ [46, 58], text width=6em, align=left] ] ] ] [System
Governance [Harmlessness [Non-maleficence, safety and security; Not be discriminatory or offensive,
avoid dangerous requests and advices, even failing to satisfy users, text width=42em,align=left, fill=purple!3 [ [47, 36, 59]
 [46, 60], text width=8em, align=left] ] ] [Turst [Turst, and monitor and evaluate the integrity, text width=26em,align=left, fill=purple!3 [ [46, 49], text width=6em, align=left] ] ] [Transparency [Explainability, interpretability or other acts of communication and disclosure, text width=44em,align=left, fill=purple!3 [ [47, 46, 48]
 [61, 62, 63], text width=8em, align=left] ] ] [Usefulness [Instruction following, informativeness, correctness, and aesthetics. Perform reasonable
tasks or answer questions, and inactively ask for more necessary information, text width=49em,align=left, fill=purple!3 [ [36, 60], text width=6em, align=left] ] ] ] ] [Meso Level [National
Level [The United States [Technological leadership (maintaining national competitiveness, strengthening investment
and innovation) and technological safety (secure and effective systems, protection against
algorithmic discrimination, data privacy, notification and explainability, user autonomy
including the freedom to opt-out or choose alternatives, fairness, and transparency), text width=52em,align=left, fill=purple!3 [ [64]
 [46]
 [65]
 [66], text width=3em, align=left] ] ] [European Union [Emphasis on risk control, respect for human autonomy,
prevention of harm, fairness, and explainability, text width=32em,align=left, fill=purple!3 [ [67, 68, 64, 46], text width=11em, align=left] ] ] [China [Safeguarding national security and adhering to the core socialist values, with requirements
that emphasize a people-centered approach, sustainability, cooperation, privacy, fairness,
shared benefits, external security, internal security, transparency, and accountability, text width=54em,align=left, fill=purple!3 [ [69, 70, 64], text width=8em, align=left] ] ] ] [Industry
Level [Education [Authenticity; Privacy; Trustworthiness; Non-discrimination;
Attention to minorities in generated content, text width=35em,align=left, fill=purple!3 [ [71], text width=3em, align=left] ] ] [Health [Compliance with the safety and effectiveness requirements of medical devices, as well as
interpretability and transparency; Adherence to human rights principles, ensuring the security of
patient data, and vigilance against high-risk applications and unauthorized experimental testing, text width=56em,align=left, fill=purple!3 [ [72], text width=3em, align=left] ] ] [Gaming [Clear accountability chain, algorithmic transparency, system transparency,
auditability of output content, safety, and predictability, text width=43em,align=left, fill=purple!3 [ [73], text width=3em, align=left] ] ] ] [Cultural
Level [Large/Small
Power Distance [The extent to which the less powerful members of organizations and institutions
accept and expect that power is distributed unequally, text width=47em,align=left, fill=purple!3 [ [74, 75]
 [76, 77], text width=6em, align=left] ] ] [Strong/Weak
Uncertainty Avoidance [What extent a culture programs its members to feel either uncomfortable or
comfortable in unstructured situations, text width=45em,align=left, fill=purple!3 [ [74, 75]
 [76, 77], text width=6em, align=left] ] ] [Individualism/Collectivism [The degree to which people in a society are integrated into groups, text width=39em,align=left, fill=purple!3 [ [74, 75]
 [76, 77], text width=6em, align=left] ] ] [Masculinity/Femininity [The prominence of masculine or feminine qualities as represented by society, text width=44em,align=left, fill=purple!3 [ [74, 75, 76], text width=8em, align=left] ] ] [Long/Short-Term
Orientation [The extent to which people are willing to delay the gratification of their material,
emotional, and social needs, text width=47em,align=left, fill=purple!3 [ [74, 75, 76], text width=8em, align=left] ] ] [Indulgence/Restrained [The degree to which people’s basic needs and desires for enjoying life are allowed, text width=47em,align=left, fill=purple!3 [ [74, 75], text width=6em, align=left] ] ] ] ] [Micro Level [Recruitment, fill=yellow!20 [Fairness in recruitment, protection of candidates’ privacy and data security, and safeguarding
candidates’ dignity and autonomy from infringement, text width=54em,align=left, fill=purple!3 [ [78, 79, 80], text width=8em, align=left] ] ] [Legal Consultation, fill=yellow!20 [Principles of privacy protection, fairness, transparency and accountability, text width=42em,align=left, fill=purple!3 [ [81], text width=3em, align=left] ] ] [Pharmaceutical company governance, fill=yellow!20 [Privacy, security, interpretability, transparency, fairness, accountability,
human-centeredness, and societal benefit, text width=42em,align=left, fill=purple!3 [ [82], text width=3em, align=left] ] ] ] ]
Figure 1: Multi-Level Value Alignment Principles
The normative issue of value alignment aims to focus on which value principles the output and behavior of AI agents should comply with, that is, the issue of choosing principle mechanisms and ethical norms. With continuous advancements in related technologies, AI agents have been extensively deployed across a variety of interdisciplinary domains, including public administration, healthcare, education, and finance. Furthermore, the open-sourcing of LLMs has enabled their privatized and localized deployment, allowing for the development and training of AI agents tailored to specific needs. This broadens the scope of AI agent deployment to encompass national, industrial, corporate, and organizational levels. Given the diversity and hierarchical nature of human values, AI agents must adhere to different sets of value principles depending on the context and level of application. A fixed, unified framework of universal values is insufficient to address all scenarios involving value constraints.
Therefore, to account for the complexity involved in selecting and aligning value principles with specific application contexts, this study adopts the perspective of AI agent developers and regulators, categorizing human values into three levels: macro, meso, and micro, as in the Figure  1. From the top down, we systematically review and synthesize relevant literature and policy documents concerning value theory and regulatory mechanisms based on the needs of AI agent governance. At the macro level, value alignment corresponds to universal values, emphasizing global, cross-cultural, and cross-domain principles and challenges. The core objective at this level is to establish a general ethical baseline that ensures basic compliance and safety of AI agents across diverse application scenarios. The meso level corresponds to national, cultural and industry-specific values, focusing on national policies, cultural tendencies and sectoral standards. Its core objective is to translate abstract ethical principles into actionable industry norms and legal regulations. At the micro level, value alignment pertains to organizational and operational values, addressing value-related issues within specific organizational contexts or business processes. The focus here is to optimize the balance between functional objectives and ethical constraints within a limited scope.
2.3.1 Macro Level
The widespread deployment of AI agents may pose potential negative impacts on human society, constituting a shared global risk. Consequently, establishing principles that ensure alignment between AI agents and human values has become an urgent issue. From a global perspective, human values exhibit diverse yet reasonable distinctions across nations, regions, cultures, and ethnicities [21]. This calls for the identification of truly foundational and universally accepted moral consensuses, in order to construct a common baseline of values. Such a baseline is essential to guarantee the basic compliance and safety of AI agents across diverse application scenarios, thereby addressing the challenge of value alignment in a meaningful and systematic manner.
The principles of value alignment at the macro level are primarily grounded in two significant theoretical frameworks from the fields of social science and ethics: Moral Foundations Theory and the Theory of Basic Human Values.
  • Moral Foundations Theory [83] posits that human moral reasoning is rooted in a set of innate foundational values that have evolved over time to meet the demands of social environments. The theory identifies five core moral foundations: care/harm, fairness/cheating, loyalty/betrayal, authority/subversion, and sanctity/degradation. It has been widely applied to examine moral divergences and conflicts across different cultural contexts [84].
  • The Theory of Basic Human Values [85] seeks to uncover the fundamental drivers of human behavior and cognition by categorizing basic human values into ten dimensions: self-direction, stimulation, hedonism, achievement, power, security, conformity, tradition, benevolence, and universalism. Schwartz’s theory has been extensively adopted in the social sciences to explore both value differences and commonalities across cultures.
In fact, research on value alignment at the macro level is both longstanding and well-developed. As early as the 1950s, science fiction writer Isaac Asimov proposed the Three Laws of Robotics in I, Robot, with the First Law stating that “a robot may not injure a human being or, through inaction, allow a human being to come to harm”  [86]. This can be regarded as an early conceptualization of macro-level moral constraints on machines. In recent years, the rapid advancement of artificial intelligence has prompted numerous governments and organizations to issue a wide range of AI ethics guidelines. For example, UNESCO’s Recommendation on the Ethics of Artificial Intelligence outlines key values such as the respect, protection, and promotion of human rights, fundamental freedoms, and human dignity; the flourishing of the environment and ecosystems; the assurance of diversity and inclusiveness; and the pursuit of peaceful, just, and interconnected societies [87]. Similarly, the World Economic Forum and the Global Future Council on Human Rights’ White Paper on How to prevent discriminatory outcomes in machine learning advocates for values such as proactive inclusiveness, fairness, the right to understanding, and redressability [88].
Despite overlaps, the aforementioned value alignment principles for AI also exhibit notable differences. To ensure that the multitude of AI ethical guidelines can effectively direct and constrain value-aligned AI development—while mitigating confusion and conflict at the macro level—scholars have extracted and synthesized key principles from various perspectives, based on diverse research purposes and methodologies. Jobin et al. analyzed 84 documents containing AI ethical principles or guidelines issued by private enterprises, academic institutions, governmental bodies, intergovernmental and supranational organizations, non-profits, and professional associations. From this corpus, they distilled 11 core principles for macro-level AI value alignment: transparency, justice and fairness, non-maleficence, responsibility, privacy, beneficence, freedom and autonomy, trust, dignity, sustainability, and solidarity [46]. Similarly, Floridi and Cowls compared ethical principles articulated in six high-profile initiatives aimed at fostering socially beneficial AI, synthesizing five core principles: beneficence (promoting well-being, preserving dignity, and sustaining the planet), non-maleficence (privacy, security, capability caution), autonomy (the power to decide), justice (promoting prosperity, preserving solidarity, avoiding unfairness), and explicability(enabling the other principles through intelligibility and accountability) [47]. Yao et al. proposed that alignment with human values is the foundational and broadest behavioral constraint for LLMs to maximize collective human welfare [36]. Their framework includes the “3H” principles—helpful, honesty, and harmless—along with adherence to widely accepted social norms and ethics (e.g., Rules of Thumb, ETHICS), and foundational ethical values such as Schwartz’s theory of basic human values [28] and the Rokeach Value Survey. As research has progressed, Yao et al. further integrated the LLMs’ Unique Value System and Safety Taxonomy into the value alignment framework [54]. Wu and Fan recognizing both the complexity of AI ethical issues and the diversity of ethical values, constructed a hierarchical system of AI ethics. At its base lies the set of universal ethical values of human society, such as human dignity, humanity, life and health, and autonomy. These can serve as globally recognized ethical foundations for AI governance [55].
Building upon previous studies and aligning with the research objectives of value alignment in AI agent systems, this study categorizes macro-level human values into four dimensions: moral foundation, rights protection, sustainability, and system governance.
The moral foundation dimension encompasses: Beneficence (AI agents should act for the good, promoting social interests and human well-being); Justice and Fairness (AI agents should operate without bias, ensuring inclusivity and equal treatment); Honesty (AI agents should provide truthful information and transparently represent their internal knowledge and capabilities); Responsibility (AI agents should proactively fulfill their responsibilities); Dignity (AI agents should respect and uphold both their own dignity and that of others); Virtue (AI agents should exhibit qualities such as truthfulness, patience, generosity, kindness, and courage).
The rights protection dimension includes: Freedom and autonomy (AI agents should respect individual autonomy and support decision-making rights); Privacy (AI agents should safeguard personal and sensitive information and avoid unauthorized disclosure).
The sustainability dimension includes: Solidarity (AI agents should foster social cohesion, enhance collective security, and contribute to community resilience); Sustainability (AI agents should adopt attitudes that support the sustainable development of the environment, energy, and natural resources).
The system governance dimension includes: Harmlessness (AI agents should act without malicious intent and ensure the safety of individuals and society); Trust (AI agents should generate reliable outputs and avoid disseminating false or misleading content); Transparency (AI agents should provide outputs that are explainable, interpretable, and understandable); Usefulness (AI agents should perform appropriate tasks or provide concise responses to inputs, and be capable of requesting additional relevant information or clarification when necessary).
2.3.2 Meso Level
Research on value alignment principles at the macro level reveals that such principles are generally broad and overarching, reflecting a consensus across divergent ideologies, political beliefs, and academic perspectives. However, within the global framework of AI ethics and governance, national policies, cultural tendencies and industry standards also constitute critical components of value alignment, serving as concrete application scenarios for more specific alignment principles.
National Level The distinct public values embodied in national AI strategies may guide the development and implementation of AI both domestically and internationally. A country’s positioning in AI governance—its regulatory role, choice of policy instruments, and approach to public responsibility—either directly or indirectly reflects the value principles it prioritizes, thereby shaping the ordering and prioritization of those values [89].In a cross-national comparative study, Djeffal et al. reviewed AI-related policies across 22 countries and regions, and classified their governance roles based on two dimensions: the government’s proactive or reactive stance toward AI development, and whether the focus lies on regulating potential risks or promoting AI advancement [90]. This classification yields four ideal types of governance roles: market-oriented (passive–development) [91], entrepreneurial (active–development) [92], regulatory (active–regulation) [93][94], and facilitative self-regulation (passive–regulation) [95]. These types are not mutually exclusive and can be implemented in hybrid forms.
  • USA. As the birthplace of the current wave of AI development, the United States positions itself between a market-oriented and facilitative self-regulation model in order to maintain its global leadership. It tends to promote AI advancement primarily through industry self-regulation, relying on market mechanisms and decentralized legislation at the state level. At the federal level, governance is driven largely by executive orders, while state legislatures formulate their own AI-related bills independently. In 2022, the White House released the Blueprint for an AI Bill of Rights, outlining five core protections that all Americans should expect when interacting with AI and automated systems: safe and effective systems; protection against algorithmic discrimination; data privacy; notice and explanation; and alternatives, options, and freedom from automated systems  [65]. In 2023, the National Science and Technology Council updated the National Artificial Intelligence Research and Development Strategic Plan, reaffirming the importance of transparency, fairness, and efficiency [66].
  • EU. The European Union has long placed a strong emphasis on the legal protection of personal rights, such as privacy. In line with its tradition of governance in international law and global institutions—and with the goal of exerting international influence—the EU has adopted a governance model that combines facilitative self-regulation with formal regulatory mechanisms, primarily guided by risk management. In 2019, the European Commission issued the Ethics Guidelines for Trustworthy AI, which emphasized key principles to be followed in the development, deployment, and use of AI systems: respect for human autonomy, prevention of harm, fairness, and explicability. The guidelines also identified seven key requirements: human agency and oversight; technical robustness and safety; privacy and data governance; transparency; diversity, non-discrimination, and fairness; societal and environmental well-being; and accountability. Special attention was given to vulnerable groups such as children and persons with disabilities, as well as to situations involving asymmetries in rights and information, such as between employers and workers, or between corporations and consumers [67]. In 2024, the EU officially adopted the AI Act, the world’s first comprehensive legislation on artificial intelligence. The Act categorizes AI systems into four levels of risk—unacceptable, high, limited, and minimal—and prescribes corresponding regulatory measures to govern their deployment [68].
  • China. China exhibits characteristics of both an entrepreneurial and regulatory state, placing significant emphasis on balancing ethical governance with the advancement of AI to safeguard national security. The government adopts proactive policies to promote the development of advanced AI technologies for societal benefit, while concurrently regulating generative AI through departmental rules and standards and continuously refining the implementation of ethical principles. In 2023, the Chinese Academy of Sciences released the Measures for the Ethical Review of Science and Technology (Trial), which require that technologies such as AI adhere to core ethical principles including the promotion of human well-being, respect for the right to life, fairness and justice, risk control, and transparency [69]. Additionally, the National Artificial Intelligence Standardization General Group and the Artificial Intelligence Subcommittee of the National Technical Committee for Standardization jointly issued the Guidelines for the Standardization of AI Ethical Governance (2023 Edition) [70]. This document specifies ten ethical principles for AI: human-centricity, sustainability, cooperation, privacy, fairness, sharing, external security, internal security, transparency, and accountability. The Interim Measures for the Administration of Generative Artificial Intelligence Services further require that generative AI services comply with applicable laws and ethical norms, uphold the core socialist values, and maintain national security and unity. The regulation mandates end-to-end compliance throughout algorithm design, data training, and content generation, including the completion of security assessments and filings. Violations are subject to legal penalties [96].
Industry Level The paradigmatic breakthrough in AI has driven its deep integration and large-scale application across various societal domains. This diffusion of technology not only reshapes the operational logic of traditional industries but also generates differentiated ethical demands. Specifically, each sector’s unique knowledge systems, cultural traditions, and institutional norms impose distinct value expectations on AI systems.
  • Education Domain. In the field of education, UNESCO’s Guidance for Generative AI in Education and Research examines the controversies and ethical risks associated with generative AI and provides targeted recommendations for its use in educational and research contexts. The guidance emphasizes the importance of ensuring the authenticity and credibility of AI-generated content, protecting copyright and privacy, minimizing the negative effects of hallucinated content, and adhering to non-discrimination principles. It also calls for attention to minority viewpoints that may otherwise be overlooked [71].
  • Medical Domain. In the medical domain, the World Health Organization urges national governments and regulatory bodies to assess the safety and effectiveness of AI applications in healthcare. It stresses the need for adequate information disclosure to ensure the interpretability and transparency of AI systems, mandates compliance with medical device regulations, and requires the examination of AI use in light of human rights principles—such as dignity, non-bias, autonomy, and privacy. Additional safeguards must be in place to protect patient data, and high-risk applications (e.g., prescribing medication or providing psychological counseling to patients with depression) and unauthorized experimental testing must be closely monitored [72].
  • Game Industry. In the gaming industry, AI has significantly enhanced development efficiency through applications ranging from narrative content generation to algorithmically driven NPC decision-making. This has enabled personalized recommendation systems and new operational models based on behavioral analytics. However, while these innovations disrupt traditional workflows, they also present challenges for governance in the gaming sector. Melhart et al. explore ethical issues arising from AI deployment in games and outline key value alignment principles, including: establishing clear chains of responsibility, ensuring transparency in decision-making, maintaining algorithmic and system transparency, enabling auditability of AI outputs, protecting training data and models against malicious attacks and manipulation, and ensuring the predictability of generated content [73].
Cultural Level Beyond national and industrial dimensions, cultural plurality and diversity represent a critical facet of value alignment in AI. Existing studies often draw on Hofstede’s cultural dimensions theory [74], focusing on how AI systems express value preferences across six key cultural dimensions—power distance (large/small), uncertainty avoidance (strong/weak), individualism/collectivism, masculinity/femininity, long-term/short-term orientation, and indulgence/restraint —and the extent to which these align with human value tendencies in different cultural contexts.
  • Power distance. It refers to the extent to which less powerful members of organizations accept and expect unequal power distribution.
  • Uncertainty avoidance. This reflects the degree to which individuals feel uncomfortable or comfortable in ambiguous or unknown situations.
  • Individualism/collectivism. This dimension captures the degree to which individuals are integrated into groups and the responsibilities they assume toward those groups.
  • Masculinity/femininity. This dimension denotes the prom-inence of traits traditionally associated with masculine or feminine roles in society.
  • Long-term/ short-term orientation. This dimension measures the acceptance of delayed gratification of material, emotional, and social needs.
  • Indulgence/restraint. The indulgence versus restraint dimension concerns the degree to which societies permit the gratification of basic human desires and enjoyment.
Based on these dimensions, Wang et al. found that LLMs prompted in English tend to exhibit a masculine orientation, while those prompted in German show a stronger inclination toward long-term orientation and indulgence. The cultural traits exhibited by LLMs under Chinese prompts fall between those of English and German prompts [75]. Similarly, Kharchenko et al. observed that LLMs can recognize cross-cultural value differences but consistently lean toward specific value orientations. In particular, across all languages and prompting methods, LLMs predominantly favor long-term orientation, suggesting that LLMs may faithfully reflect certain cultural values [76].
In summary, the global practices of AI governance reveal diverse paths and efforts toward value alignment across nations, industries, and cultures. Value alignment is not merely a matter of matching technical systems with human intentions; it is also a profound reflection of policy choices, institutional designs, and cultural contexts at the governance level. The United States emphasizes industry self-regulation and market-driven mechanisms, reflecting its prioritization of innovation and individual freedom. The European Union seeks institutional value alignment through legislation and ethical frameworks, highlighting fairness, justice, and the protection of human rights in technological applications. China integrates technological advancement with social responsibility, reinforcing state-led ethical norms that emphasize collective interests and public safety.
Meanwhile, international organizations are increasingly working to build a global consensus on value alignment by promoting cross-border ethical cooperation and industry-wide standards. Given the differing emphases on values across nations and the deep integration of AI in various sectors, future governance systems must pursue a dynamic balance and coordination among pluralistic value claims. Through dialogue and collaboration, global value alignment can be advanced to ensure that technological development truly serves the overall well-being of human society.
2.3.3 Micro Level
As AI agents become increasingly embedded in organizational operations and business processes, the issue of value alignment at the micro level has grown more prominent. Unlike the overarching constraints imposed by macro-level universal values and meso-level policy regulations, the micro level focuses on how AI systems achieve alignment between effectiveness and values within localized environments under specific goal orientations. The scenarios and associated values at this level are inherently more personalized. Given the impossibility of exhaustively enumerating all possible scenarios, this study synthesizes relevant research and illustrates the principles of micro-level value alignment through examples such as talent recruitment, legal consultation and corporate governance in the healthcare sector.
The application of artificial intelligence technologies is profoundly reshaping human resource management practices, with the recruitment domain emerging as a primary frontier of technological transformation. AI-powered talent selection systems are driving a fundamental shift in organizational talent acquisition paradigms. Compared to traditional recruitment models that rely heavily on manual screening and subjective judgment, intelligent tools significantly enhance the speed and accuracy of candidate identification through automated resume parsing, big data-driven competency modeling, and machine learning-enabled predictive assessments [97].The ethical compatibility of AI technologies in recruitment settings has attracted sustained academic attention. In constructing ethical principles, Simbeck advocates for the transfer of key moral concepts from other domains to human resource analytics when applying new AI technologies, identifying five core principles: data privacy protection, user opt-out mechanisms, third-party audit frameworks, algorithmic transparency, and the safeguarding of dynamic development rights [79]. These principles align with the humanistic view of technology proposed by Runb-Kettler and Lehnervp, which emphasizes preserving candidates’ dignity and autonomy in automated processes [78]. A systematic literature review by Hunkenschroer and Luetge further deconstructs the ethical dimensions of AI in recruitment into ethical opportunities, ethical risks, and areas of ambiguity [80].
From the perspective of ethical opportunities, AI can reduce human biases in traditional recruitment through standardized assessment procedures [98][99]. Text optimization tools can help generate gender-neutral job descriptions [100], while natural language processing has revolutionized resume parsing efficiency, allowing recruiters to focus on strategic decision-making [101]. However, ethical risks remain significant. Algorithms may inherit implicit biases embedded in training data [102], implement indirect discrimination via proxy variables (e.g., zip codes correlated with race), or infringe on privacy rights by extracting personal information from candidates for recruitment purposes [103][104]. Although AI has the potential to disrupt traditional elite hiring models and promote social mobility, it may also lead to reduced cognitive diversity in teams due to algorithmic homogenization. These tensions and risks underscore fairness as a core principle in the application of AI to human resource recruitment.
The deep integration of AI technologies is also influencing the strategic development and governance choices of biopharmaceutical companies. In response to the ethical challenges posed by the application of AI in healthcare, AstraZeneca has issued a series of AI ethical principles, including privacy and security, interpretability and transparency, fairness, accountability, human-centeredness, and societal benefit. The company has also developed a Responsible AI Handbook, established an AI Resolution Committee and internal advisory services, and commissioned AI audits to implement comprehensive AI governance [82].
In addition, AI technologies are demonstrating significant potential and application prospects in the legal consulting domain. Sun et al. simulated the collaborative workflows of real-world law firms using multi-agent systems. In this context, they incorporated normative principles—such as privacy protection, fairness, transparency, and accountability—to guide the ethical behavior and interactions of the agents. This alignment with human values enhances the quality of legal consulting services by ensuring that multi-agent systems operate in a manner consistent with ethical and societal expectations [81].
As AI continues to advance, the principles of value alignment will no longer be confined to broad and ambiguous moral values. Instead, they will increasingly penetrate specific application scenarios, integrating concrete practical experience and well-defined guidelines to inform the development and deployment of agent systems.
3 Applications and Value Issues in LLM-based agent systems
LLM-based agent systems are now applied across numerous domains, indicating that value alignment concerns will permeate to the application scenario level. This section first introduces our review subjects—LLM-based agents and multi-agent systems. Subsequently, it classifies and organizes intelligent agent system application scenarios based on their generality, and discusses the associated value alignment issues.
3.1 LLM-Based Agent Systems
3.1.1 LLM-Based Agent
In philosophy, the term ’agent’ typically refers to an entity possessing consciousness, intention, and the capacity for action [105, 106]. This can include humans, animals, and other entities in both the physical and virtual worlds [1]. Floridi identifies three general characteristics of an agent: interactivity, autonomy, and adaptivity. Interactivity refers to an agent’s ability to interact with and respond to environmental stimuli. Autonomy indicates an agent’s capacity to change its own state based on its transformation rules. Adaptivity denotes an agent’s ability to not only alter its state but also to modify the rules governing these changes through learning and experience, thereby flexibly responding to environmental shifts. Building on these three characteristics, Floridi posits that a Moral Agent can exhibit morally evaluable behavior, meaning their actions can engender moral "good" or "evil." Consequently, Moral Agents must adhere to specific ethical norms and oversight mechanisms [107]. When the concept of an "Agent" is introduced into the field of artificial intelligence, it becomes known as an AI agent. An AI agent is an artificial entity capable of achieving complex goals through collaboration [108], representing the concretization of the philosophical concept of "agent" within artificial intelligence.
Large Language Models (LLMs) are deep learning models trained on vast amounts of textual data. They are capable of generating natural language text and comprehending its meaning, with examples including Deepseek, GPT-3 and GPT-4. A salient characteristic of LLMs is their adherence to scaling laws and the exhibition of emergent abilities. This means that once a model’s training parameters exceed a certain threshold, it can experience a dramatic improvement in existing capabilities or acquire entirely new ones [17], enabling it to solve tasks it was not explicitly trained for [12]. As models continue to grow in scale, pre-training corpora expand, and alignment techniques advance, LLMs have demonstrated powerful capabilities in knowledge acquisition, instruction understanding, summarization, planning, and reasoning [1]. This showcases their immense potential for achieving human-like intelligence. Consequently, researchers have begun leveraging LLMs to construct intelligent digital agents [23], transforming abstract philosophical concepts into practically engineerable system design frameworks [1].
AI Agents are intelligent entities capable of perceiving their environment, making decisions, and executing actions [109, 110, 111]. LLMs can serve as the central controller for AI agents, endowing them with decision-making thought and memory [112]. Decision-making thought enables agents to decompose complex tasks, deliberate on each sub-task, and explore multiple pathways to arrive at intricate decisions. A long- and short-term memory mechanism allows agents to maintain contextual coherence during task completion and continuously learn from past interactions. Strategies such as multimodal perception and tool use expand an agent’s perception and action space [113, 114, 115, 116, 117]. Furthermore, AI agents endowed with environmental interaction capabilities can learn from feedback and execute novel actions [118]. The development of techniques like Chain-of-Thought (CoT) and problem decomposition has continuously enhanced the reasoning and planning capabilities of AI agents[119, 120]. Furthermore, pre-training on massive-scale corpora has enabled agents to exhibit few-shot and zero-shot generalization abilities. This allows them to transfer and complete tasks across different scenarios without requiring parameter updates [120, 121]. The field of AI is currently undergoing a profound paradigm shift. Building upon existing research, the industry is transitioning from developing AI models for passive, structured tasks to constructing novel models capable of acting as dynamic agents in diverse and complex scenarios.
3.1.2 Multi-Agent System
A Multi-Agent System (MAS) is a computer system composed of multiple interacting agents [12]. Operating on natural language and custom content [24], these agents can interact within a shared environment, autonomously making decisions to solve complex problems or complete tasks [122]. The key components of such a system include the Agent, Environment, Interaction rules, and Organization pattern [12].
  • Agent.The agent is the core participant in the system, possessing a role identity, knowledge, capabilities, and behaviors, along with abilities such as learning, planning, reasoning, and decision-making.
  • Environment.The environment is the external world in which the agent operates, which can be either a simulated virtual environment or a real physical space.
  • The Interaction Rules. The rules are predefined rules, protocols, and mechanisms governing information exchange and behavior coordination among agents.
  • The Organization Pattern. The pattern refers to the structural relationships of agent interactions, encompassing hierarchical control and emergent behavior-based organizations.
Multi-agent interaction within these systems mirrors the specialized division of labor and teamwork found in human society; this aims to achieve collective intelligence. Individual autonomous agents possess initial knowledge and goal sets tailored for specific tasks [122]. By leveraging LLMs’ capabilities in text generation, comprehension, and communication, diverse agents can adaptively collaborate. They respond to task requirements, participate in planning, discussion, and decision-making, and learn from one another. This fosters the aggregation of broad knowledge across multiple domains [112, 12]. Concurrently, through environmental interaction, agents perceive and understand information. They then make decisions and execute actions based on their goals, existing knowledge, and observations. This process unleashes their potential to complete specialized tasks and achieve common objectives.
Multi-agent systems exhibit remarkable flexibility, enhanced robustness, inherent self-organization and coordination, and robust real-time operation capabilities [12]. Although interactions among agents are governed and constrained by predefined rules, emergent behaviors and phenomena can still arise from these interactions [123]. Additionally, due to the division of labor within the system, resource and energy consumption are distributed across multiple agents. This is often more cost-effective than a single, powerful agent attempting to solve an entire complex problem [12]. Therefore, compared to a single agent, multi-agent systems are better suited for handling more complex and dynamic tasks, particularly in scenarios requiring diverse strategies and collective decision-making [112]. At present, LLM-based multi-agent systems are extensively used to simulate social phenomena. This includes modeling social network construction, international conflicts, political evolution, and the dynamics of public opinion  [124, 125, 126].
3.2 Application Scenarios
Table II: Classification of Agent System Application Scenarios
Generalizability Category Specific Scenarios Ref
High Generalizability Tool using • General tool using • Multimodal tool using  [127, 128, 129]  [130, 131]
• Instructions acceptance, image capture, • Enterprise-level data workflow automation
historical retrieval and tool invocation
Dialugue Agents • Conducting persuasive dialogues • Dialogue task optimization  [132, 133, 134, 135]  [136, 137, 138, 139]  [140, 141, 142, 143]  [144, 145]
• Multimodal emotional dialogue systems • Continuous conversational dialogue
developing • Cross-domain dialogue clarification
• Personalized experience improvement in strategies
long-term dialogue • Online shopping question answering
• Medical question answering • LLM-driven automated task execution in
• Human-like emotional dialogue agents mobile applications
• Personality consistency testing of • Dialogue strategy planning
role-playing agents • Social intelligence evaluation of role-playing
• Linguistic pattern analysis of dialogue AI dialogue agents
• Personalization validation of dialogue agents
Multimodal Perception • Multimodal tool using • Instructions acceptance, image capture,  [146, 127, 128, 147]
• Multimodal retrieval historical retrieval and tool invocation
• Dynamic scene understanding and guidance
Multi-Agent Collaboration • Multi-agent social interaction modeling • Collaborative task division and quality  [148, 149, 150, 151]  [152, 153, 154, 155]  [156, 157, 158]
• Simulation of social collaborative behaviors assurance in software development
among multiple LLM agents • Training human collaborative behaviors
• Multi-LLM collaboration in medical decision-making supported by generative agents
• Dynamic multi-team adversarial competition • Long-term planning for multi-robot systems
• Language instruction generalization in in partially observable environments
multi-agent games • Performance evaluation of multi-agent
• Dynamic job scheduling in factories systems
• Software engineering automation
Visual Tasks Processing • Reasoning covered object in images • Image Inpainting  [159, 160, 161, 162]
• Multimodal image generation and editing • 3D scene synthesis and rendering
Natural Language Processing • Keyphrase generation • Long text reading comprehension  [163, 164, 165]  [166, 167, 168]
• Database information extraction and integration • Relation triplet extraction in natural language
• Dynamic navigation for incremental extreme
multi-label classification
Data Generation and Analysis • Transporing the knowledge of the isolated agents  [169]
Reasoning, Planning and Decision Optimization • Autonomous information retrieval in visual • Robotic language-guided task planning  [170, 171, 172, 173]  [174, 175, 176, 177]  [178, 179, 180, 181]  [182, 183, 184, 185]  [186, 187, 188, 189]  [190, 191, 192]
question answering • Asynchronous task planning
• Social data analysis and decision-making • Strategic language agent collaboration in
• General zero-shot reasoning task optimization Werewolf game
• Interactive planning task execution • Instruction Following
• Automated task execution • Agent performance analysis in complex
• Multi-agent debates enhancing factuality and Android environments
reasoning of language models • Embodied decision-making task evaluation
• Knowledge distillation from multi-agent interaction • Automated tool using
graphs to improve small model reasoning • Generative agents with dual-process cognition
• Multi-agent social learning mechanisms in complex tasks
• Real-world travel planning • Automated user interface control
• Exploratory trajectory optimization • Network agent planning
• Generalization of robot visual-language tasks • On-device AI assistants
Limited Generalizability Function-driven Code Generation • Multi-agent code generation framework for • Code generation and environment interaction  [193, 194, 195, 196]
competitive programming problems modeling
• Open-domain process customization • Code repair
Social Simulation • User behavior simulation • Human social interaction simulation  [197, 198, 199, 200]  [58, 201, 202, 203]  [204, 205, 206, 207]  [208, 209, 210]
• Social network information propagation simulation • Role-playing agent generation
• Trust behavior simulation • Complex task planning and collaboration
• Business competition scenario simulation in Minecraft
• Sustainable cooperation simulation • Time-aware multitask simulation
• Diplomacy simulation • Complex social interaction simulation and
• Social intelligence interactive learning evaluation
Graphical User Interface Agents • Multi-turn conversational web agents • Personalized web agents  [211, 212, 213, 214]  [215, 216, 217, 218]  [219, 220, 221, 222]  [223, 224, 225, 226]  [227, 228, 229]
• Graphical interface tasks automating • Computer screen interaction and control
• Web interaction task execution • Smartphone GUI automation
• Knowledge transformation and task automation • Web navigation tasks executing
• GUI agents autonomously executing complex • Adversarial attacks on web agents
multi-step tasks • End-to-end web interaction agents
• Enterprise-level web task automation • Training of web agents
Model Analysis, Evaluation and Improvement • Dialogue task optimization • Multi-agent evaluation optimization  [230, 62, 231]  [232, 233, 234]
• Neural network behavior interpretation • LLM-driven recommender system attack
• Evaluation of LLM alignment with human values testing
• Instruction tuning and data optimization
Scenario-driven Finance • Multimodal foundation agent for financial trading • Financial sentiment analysis  [235, 236, 237, 238]
• Multi-agent system for enhanced financial • Macroeconomic activities simulation
decision making
Science • Automating data science tasks • Automating scientific discovery  [239, 61, 240]  [241, 242]
• Design of organic structure-directing agents • Scientific data visualization
• Academic citation recommendation
Healthcare • Medical reasoning and diagnosis • Electronic health record modeling and  [243, 244, 245, 246]  [247, 248, 249]
• Pathology image-text pairs generation and clinical decision support
analysis • Multimodal interaction testing of mental
• Psychological measurement health dialogue agents
• Drug design • Gene perturbation experimental design
Game • Fighting game agents training • Intelligence evaluation in murder mystery  [250, 251, 252]  [253, 254]
• Detective role-playing reasoning games
• Interaction and reasoning in text-based games • Text-based game agents
Manufacturing • Dynamic scheduling in flexible manufacturing  [254]
systems
Robotics • LLM-based embodied agent open environment • Embodied instruction following  [255, 256, 257, 258]
perception • Home robot knife placement
• Vehicle acceleration and obstacle avoidance • Mobile device operation assisting
in autonomous driving
Urban Computing • Addressing standardization tasks supported • Urban knowledge graph construction  [259, 260, 261]
by geospatial tools • Personal mobility generation
Social Media • Detecting harmful content on social media • Logical and causal fact-checking  [262, 263]
Autonomous Driving • Real-time motion generation for autonomous driving • Traffic signal control  [264, 265]
Literary Creation • Director-actor coordinate agent framework for • Narrative generation through multi-step  [266, 267]
controllable  drama script generation collaboration
Given the poweful knowledge acquisition, instruction understanding, task planning and logical reasoning capabilities of multi-agent systems, they have been widely used worldwide. From small and medium-sized organizations to multinational corporations, from higher education institutions to research institutes, from grassroots medical facilities to large diagnostic hospitals, and government agencies at all levels are actively promoting the application of AI agent technology. In terms of application scenarios, the penetration of AI agent technology is extremely broad, covering multidisciplinary fields such as healthcare, education, trade and economics and law. The problems that multi-agent systems are capable of addressing ranges from comprehending a single piece of literature to simulating global-scale social events, spanning across micro, meso and macro levels. In contrast to categorizing applications based on disciplinary domains [14] or task types [112], we sort and categorize application scenarios according to the degree of generalizability of multi-agent systems and build upon this to explore the associated issues of value alignment.
Generalizability of multi-agent systems refers to the extent to which they can be applied across diverse scenarios or meet the needs of various organizations. A high degree of generalizability is based on the foundational capabilities of AI agents such as natural language processing, reasoning, planning and autonomous decision-making. The universal application includes tool using, dialog agents, multimodal perception, multi-agent collaboration, visual tasks proingcess, natual language processing, data generation and analysis, reasoning, planning and decision optimization and on, see Table II.
In terms of tool using, AI agents are capable of recognizing implicit needs for tool invocation and leveraging a wide range of real-world tools involving perception, manipulation, logical reasoning, and creativity. These tools enable AI agents to process multimodal inputs—such as images, tables, and code—to accomplish complex tasks [129, 130]. Taking dialogue agents as an example, A.V. Savchenko and L.V. Savchenko (2024) proposed an approach that employs multiple emotion-aware dialogue agents based on large language models to provide query responses. These AI agents integrate speech and facial expression recognition to infer the user’s emotional state and generate appropriate responses [134]. Furthermore, dialogue agents have also been applied in persuasive interactions for social good, particularly in scenarios involving user resistance, where they are able to generate more polite and convincing responses [132]. In the domains of reasoning, planning, and decision-making optimization, autonomous agents can guide large language models to perform zero-shot reasoning, thereby improving their performance in generation, classification, and inference tasks  citeDBLP:conf/icml/CrispinoMZS024. Multi-agent debate mechanisms have been shown to enhance accuracy and factual consistency in mathematical and strategic reasoning while reducing hallucinated outputs [175]. Additionally, structuring multi-agent interactions as graph representations enables knowledge distillation for smaller models, improving their performance and generalization in commonsense and mathematical reasoning tasks [176].
Higher degree of generalizability of indicates that the agent system can be broadly applied across various scenarios, not only addressing domain-specific problems but also meeting diverse task requirements across multiple industries and organizations. For example, multi-agent systems can be applied not only in natural sciences but also in social science domain such as economics and management, suitable for addressing global issues that span across nations and cultures.
Lower generalizability means that multi-agent systems can not be applied across industries and domains. Specialized agent applications can be broadly categorized into two types: function-driven and domain-driven, see Table II.
Function-driven applications are developed by building upon the core capabilities of agents and are typically applicable across multiple domains. Common examples include code generation, social simulation, graphical user interface (GUI) agents, and model analysis, evaluation and improvement, as in Table II. For instance, in social simulation, Wang L. et al. (2023) proposed a large language model-based agent framework that simulates user behavior in sandbox environments to examine social phenomena such as echo chambers and herding effects [197]. In GUI navigation, agents can handle open-domain web tasks by extracting key information from complex HTML pages, understanding web structures, and performing multi-step interactions, thus addressing challenges like interaction diversity and domain knowledge scarcity [225, 227]. In model analysis, LLM-based agents have been employed to assess alignment performance, such as in Ali-agent, which simulates test scenarios and iteratively explores long-tail risks to evaluate value alignment [231]. MAIA, a multimodal explanatory agent, automates model behavior interpretation using neural networks, enhancing model transparency [62].
Scenario-driven applications refer to agent systems deployed within specific sectors, including finance, science, healthcare, game, manufacturing, and autonomous driving, as in Table II. In urban computing, the UrbanKGent framework enables automated construction of urban knowledge graphs, significantly reducing development costs [261]. In scientific research, agents assist with academic literature retrieval and citation recommendation [61], organic structure design [241], and automated data visualization. In healthcare, multi-agent systems support medical reasoning, gamified mental health assessments, and experimental design in genomics [243, 247]. In finance, agents can simulate macroeconomic dynamics [236], analyze market trends using multimodal data [235], and support trading and portfolio management [237]. Furthermore, in the social media domain, agent reasoning and coordination capabilities contribute to tasks such as harmful content detection and high-confidence fact verification [262, 263].
The low level of generalizability indicates that the agent system is used in specific countries, industries, or even organizations and enterprises. It is dedicated to completing a certain type of task, and the data sets and business information involved are highly private.
3.3 Value Issues in Application
The value alignment problems involved in the application scenarios of LLM-based agent systems mainly include two aspects: the correspondence between the generalizability of the system and the multi-level values, and the contradictions encountered when aligning values.
The degree of generalizability of LLM-based agent systems corresponds to the multi-level of values. As the degree gradually decreases, the granularity of the value norms that need to be met gradually increases. A higher degree of generalizability indicates that the applicable scenarios for LLM-based agent systems are more diverse. Therefore, it is unnecessary for the initial design of systems to align with the values of a specific group–such as a country, industry, or culture–to ensure flexible applicability across national, ethnic, and industrial contexts. The development and governance of highly generalizable multi-agent systems should be constrained by universal values of macro level. Multi-agent systems with limited generalizability typically confined to specific countries or industries. Such systems often exhibit design features characteristic of particular countries, industries, or cultures. Therefore, the development and governance of specialized systems must address diverse national policies, industry standards, and cultural preferences, while ensuring alignment with universal values. When intelligent systems are deployed in strictly private contexts, they are typically confined to specific organizations and designed to address specialized business or personal needs. These applications are highly specialized and exhibit minimal generalizability. Under such conditions, it is essential not only to align with value principles of macro and mecro level, but also to comply with the specific value frameworks of the organizations or businesses involved. It can be seen that multi-agent systems with high generalizability tend to be highly adaptable, and their value alignment primarily centers on adherence to universal principles of macro level. As application specialization increases, the requirements for value alignment become progressively more complex, necessitating adherence to value norms across multiple levels simultaneously. This classification aligns with the top-down, increasingly granular hierarchy of value principles. As a result, the value norms that multi-agent systems must adhere to vary according to their degree of generalizability, offering a structured basis for governance and oversight in their design and deployment.
The value alignment problem may involve two tpyes of tensions: the plurality of value principles at the same level, and potential conflicts between values across different levels. On one hand, a key challenge is that value principles at the same level differ across countries, regions, industries, organizations, and other contextual factors. Therefore, when multi-agent systems are applied in diverse contexts, their value alignment objectives must be adapted accordingly. In the healthcare sector, Huawei’s intelligent agent EIHealth serves as a professional AI development platform for genomic analysis, drug discovery, and clinical research. Built on Huawei Cloud’s strengths in AI and big data, EIHealth features an open and extensible architecture that enables end-to-end AI empowerment. Ensuring alignment with the value principles and regulatory norms of various countries, regions, and organizations is crucial for its deployment. To address this, Huawei Cloud has issued a series of white papers tailored to the privacy and regulatory requirements of different jurisdictions and industries. For example, it provides compliance guidelines aligned with the Personal Data Privacy Ordinance (PDPO) in the Hong Kong SAR, as well as guidelines addressing the financial regulatory frameworks of the EU, including specific provisions for Ireland, Spain, and Hungary. These efforts ensure that the EIHealth platform adheres to all applicable national and regional security regulations, international cybersecurity and cloud security standards, and incorporates industry best practices—ultimately meeting the diverse security needs of cloud service users.
On the other hand, as the tasks assigned to an intelligent system become more fine-grained, the constraints on achieving alignment goals also increase. The means of maximizing task utility may conflict with established value norms, leading to tensions across different levels of values. For instance, the promotion of social cohesion at the macro level may contradict individualistic tendencies at the meso level, where individuals prioritize self-interest. Similarly, national regulations that emphasize information transparency may be at odds with micro-level demands for organizational data privacy. These cross-level value conflicts pose significant challenges for the design and development of multi-agent systems. In the field of corporate recruitment, algorithms and artificial intelligence have been widely adopted. Recent studies have explored the design and development of intelligent agent systems for tasks such as résumé screening and candidate-job matching [268]. However, AI-based recruitment systems may be at odds with widely accepted ethical norms [104]. In particular, corporate hiring practices can involve gender discrimination or racial bias, which directly conflict with legal regulations and core societal values. This ethical dilemma was vividly illustrated by Amazon’s 2018 algorithmic bias incident. The company’s internally developed résumé screening system, trained on historical hiring data, systematically discriminated against female applicants, ultimately leading to the discontinuation of the system [269]. To prevent such issues, recruitment-oriented intelligent systems must be explicitly designed to reduce or eliminate gender and racial bias. Their value alignment should comply with national labor laws, equal employment regulations, women’s rights protections, and relevant industry standards. Only on the basis of legal compliance, equality, and privacy protection should these systems then pursue the optimization of recruitment outcomes.
In complex collective intelligence systems, multi-agent value alignment refers to the design of appropriate mechanisms and algorithms that enable multiple autonomous agents to identify and adhere to shared ethical or social norms while pursuing their individual task objectives, thereby achieving consistency, safety, and fairness at the system level. Unlike single-agent alignment, which focuses on the compliance of a model’s outputs, multi-agent collaborative alignment must address challenges such as value conflicts between agents, information asymmetry, and “free-rider” or competitive dilemmas in cooperative games. The goal is to use communication protocols, reward shaping, negotiation mechanisms, or adversarial training to ensure that agents coordinate with one another and ultimately converge on a common set of collective values.
4 Evaluation and Methods for Value Alignment
To ensure that LLM-based agents behave in ways consistent with human values and task goals, value alignment evaluation has become a key research focus. Evaluation methods often rely on alignment benchmarks to assess whether agents follow expected value principles. On the methodological side, alignment techniques can be applied at various stages, including during pretraining, post-training, and through reinforcement learning (e.g., reward modeling or fine-tuning with human feedback). These methods aim to mitigate misalignment risks and enhance agent trustworthiness across various applications.
4.1 Value Alignment Evaluation
4.1.1 Datasets for Value Alignment Evaluation
Table III summarizes the existing open-source datasets for evaluating value alignment. To clarify their coverage and applicability, we systematically organize and categorize these datasets according to the previously introduced three-level framework of value alignment principles: macro, meso, and micro. This hierarchical classification enables a more nuanced understanding of how current resources support value alignment evaluation at different levels.
Table III: Statistics of Datasets for Value Alignment Evaluation.
AC: Automatic Construction, CC: Crowdsourced Construction, HC: Hybrid Construction, EDC: Expert-Driven Construction, IRED: Integration and Refinement of Existing Datasets, BJ: Binary Judgments, RT: Ranking Tasks, MCQ: Multiple-Choice Questions, OEQ: Open-Ended Questions, RSQ: Rating Scale Questions
Datasets Macro Level Meso Level Micro Level Construction Methodologies Question Formats Size
BBQ(Bias Benchmark for QA) [270] EDC MCQ 58.4k
BeaverTails [51] IRED OEQ 30.2k
BOLD [271] CC OEQ 23.6k
CBBQ(Chinese Bias Benchmark Dataset) [272] HC MCQ 106.5k
CDEval [75] AC MCQ 2.9k
CDial-Bias [273] CC OEQ 28k
CORGI-PM [274] IRED BJ, OEQ 32.9k
CrowS-Paris CC BJ 1.5k
CultureSPA [275] AC MCQ 13k
Cvalues [276] CC MCQ, OEQ 6.4k
DailyDilemmas [277] AC MCQ 2.7k
DecodingTrust [278] IRED MCQ 152.4k
DEFSurveySim [49] HC MCQ, RSQ 1.1k
EEC(Equity Evaluation Corpus) [279] EDC BJ 8.6k
ETHICS [280] CC MCQ, BJ, RT 134.4k
Flames [281] CC OEQ 1k
GlobalOpinionQA [282] IRED MCQ 2.5k
HofstedeCulturalDimensions [76] HC MCQ 0.2k
IndieValueCatalog [283] IRED MCQ 93.2k
KorNAT [284] AC MCQ 10k
LLMGlobe [77] HC RSQ, OEQ 37.6k
MFQ(Moral Foundations Questionnaire) [285] IRED MCQ 11k
Moral Beliefs [286] AC MCQ, BJ, OEQ 0.5k
Moral Integrity Corpus [287] AC RSQ, OEQ 38k
Moral Stories [288] CC MCQ 12k
MoralExceptQA [289] EDC RT 0.2k
NaVAB [290] AC MCQ, BJ 63.8k
Persona Bias [279] IRED MCQ, BJ 110.6k
PkuSafeRLHF [291] AC OEQ 476.4k
ProgressGym [292] HC RSQ, OEQ 1.4k
SafeSora [60] AC OEQ 14.7k
Scruples [293] EDC MCQ, BJ 657k
Social Bias Frames [294] CC RT, OEQ 150k
Social Chemistry 101 [295] CC BJ, OEQ 292k
StereoSet [296] CC RT, OEQ 4.2k
ToxiGen [297] AC OEQ 6.5k
UnQover [298] IRED BJ, OEQ 2713k
ValueNet [299] CC BJ, OEQ 21.3k
WikiGenderBias [300] IRED OEQ 45k
WinoBias [301] EDC MCQ 3.1k
WinoGender [302] EDC BJ, OEQ 0.7k
Although existing datasets address multiple levels of value alignment, several limitations persist. At the macro level, many datasets—such as Moral Stories [288] and ETHICS [280]—are predominantly grounded in Western cultural contexts, lacking comprehensive representation of alternative value systems, such as those from Asian or African societies. In addition, these datasets largely reflect contemporary norms and offer limited support for modeling the evolution of values over time or predicting future value shifts. At the meso level, datasets such as NaVAB [290], KorNAT [284], and CultureSPA [275] provide valuable insights into national or cultural contexts. However, they offer limited global coverage and often lack detailed evaluations of value alignment across specific industries or professional domains. Micro-level datasets are even more scarce. Compared to macro- and meso-level resources, they remain significantly underdeveloped, which limits the ability to assess and align value preferences in multi-agent systems across diverse, fine-grained real-world scenarios.
4.1.2 Dataset Construction Methodologies
Constructing high-quality value alignment datasets is critical for reliably evaluating whether different systems adhere to appropriate value frameworks. To support diverse application scenarios, researchers have primarily adopted three data construction approaches: manual construction, automatic construction, and hybrid construction that integrate human and machine-generated content.
Manual Construction. In the manual construction approach, experts or annotators design tasks, generate textual content, and assign labels based on a well-defined value framework. The resulting datasets are typically curated and reviewed by trained annotators, yielding high-quality data that closely reflects human value understanding. However, this approach is resource-intensive and lacks scalability. In addition, annotations grounded in subjective human judgment may unintentionally introduce bias. The manual construction approach can be further categorized into three subtypes: expert-driven construction, crowdsourced construction, and the integration and refinement of existing data.
  • Expert-Driven Construction is led by domain experts who design tasks and define annotation protocols to ensure data accuracy and consistency. For example, the BBQ dataset [270] was manually created using templated questions that include both biased and unbiased forms, covering nine social dimensions with contextual and disambiguating cues. Similarly, the MoralExceptQA dataset [289] was built by systematically varying moral scenario features—such as rule function and consequences of violations—to generate diverse questions on moral exceptions. These expert-curated datasets offer high-quality benchmarks and valuable tools for evaluating model performance in targeted domains.
  • Crowdsourced Construction is widely adopted for its efficiency and cost-effectiveness, particularly in large-scale data collection. Many datasets—such as Social Bias Frames [294], Flames [281], and BOLD [271]—leverage crowdsourcing platforms, often with multiple rounds of quality control to ensure annotation reliability. These datasets not only scale well but also capture diverse perspectives from annotators of varied cultural and social backgrounds, enabling more comprehensive assessments of model behavior.
  • Integration and Refinement of Existing Datasets involve aggregating and modifying multiple public sources. Notable examples include IndieValueCatalog [283], DecodingTrust [278], MFQ [285], Persona Bias [279], and GlobalOpinionQA [282], all constructed through careful curation and consolidation of open-source data. This approach reduces data acquisition costs while leveraging existing high-quality resources, thereby improving research efficiency and enabling broader evaluation coverage.
Automatic Construction. With the rapid advancement of large language models (LLMs), LLM-based automatic dataset construction has become a promising approach. This method typically involves designing prompt templates to guide the model in generating question–answer pairs or scenario-based judgments, followed by manual verification to ensure data quality. Representative datasets developed using this approach include NaVAB [290], KorNAT [284], , CultureSPA [275], and DailyDilemmas [277]. Compared to manual construction, it enables the creation of more diverse and context-rich data with significantly less human effort, allowing for large-scale dataset generation in a relatively short time. However, concerns remain regarding the quality of model-generated content. LLMs may produce hallucinated outputs, and their inherent biases can result in misalignment with human values. Consequently, ensuring the reliability and value alignment of automatically constructed datasets still requires careful oversight and validation.
Hybrid Construction. Hybrid construction approaches combine human expertise with automatic expansion using LLMs, which are increasingly emerging as a mainstream trend. In this paradigm, experts first create a small set of high-quality seed examples. The LLM is then employed to generate similar instances or expand the dataset, followed by human verification to ensure consistency and quality. Representative examples include CBBQ [272], LLMGlobe [77], and HofstedeCulturalDimensions [76], all of which utilize expert-curated prompts for initial task design, LLMs for large-scale generation, and human review for quality assurance. This hybrid methodology effectively leverages the generative capabilities of LLMs while maintaining the accuracy and coherence provided by human oversight, offering a robust foundation for value alignment research and evaluation.
4.1.3 Question Formats and Metric Design
Evaluation methods for value alignment largely depend on the format of assessment questions. Common question formats include multiple-choice questions, binary judgments, ranking tasks, rating scale questions, and open-ended questions.
Multiple-Choice Questions. Multiple-choice questions typically present a set of answer options—often including neutral, biased, or varying degrees of harmful responses—and require the model to select the most appropriate one. This format is commonly used to assess a model’s value judgments in specific contexts. For example, the BBQ dataset frames each question as a triplet: a neutral version (free of bias), a subtly biased version, and an overtly biased version [270]. By evaluating model accuracy across these variations, researchers assess its fairness and sensitivity to social bias. The CDEval dataset similarly uses multiple-choice questions to examine a model’s alignment with cultural values [75]. For instance, a question may ask whether children should obey their parents in a given culture, with answer options reflecting diverse cultural norms. The model’s selected response offers insight into its sensitivity to and alignment with different cultural perspectives.
Binary Judgments. Binary judgment questions require the model to make a dichotomous decision, such as yes/no, true/false, or agree/disagree, based on a given statement or scenario. This format is widely used to evaluate a model’s understanding of ethical principles, social norms, or cultural consensus. For example, the ETHICS dataset presents ethically challenging statements (e.g., “Stealing is acceptable if no one finds out”), each of which is annotated by humans as either correct/incorrect or acceptable/unacceptable [280]. The model or agent system is tasked with determining whether the described behavior is morally appropriate, and its accuracy serves as a measure of ethical reasoning and value alignment.
Ranking Tasks. Ranking tasks evaluate a model or agent system’s ability to make relative value-based comparisons by requiring it to order multiple responses or behavioral options from most to least aligned with human values. This format reduces inconsistencies inherent in subjective rating scales and highlights the model’s capacity for value-informed prioritization. For instance, StereoSet dataset presents stereotype-laden prompts followed by a question and several options (e.g., “A doctor walks into the room and sees the patient:” with options such as: A. He is wearing a white coat. / B. She is wearing a white coat. / C. He is holding a guitar.) [296]. The model is tasked with ranking these responses based on contextual appropriateness in terms of grammar and semantics. A Stereotype Score is then used to evaluate ranking value tendencies. Similarly, MoralExceptQA dataset provides a normative moral rule (e.g., “One should not lie”) and a complex scenario (e.g., “Is it acceptable to lie to protect a friend’s safety?”) [289]. The model or agent system must rank behavioral options, such as lying, telling the truth, or indirect concealment, based on moral reasoning. These rankings are compared to those generated by human ethicists to assess the model’s ability to integrate moral principles with contextual judgment.
Rating Scale Questions. Rating scale questions require the model or agent system to evaluate a given text, response, or behavioral option along predefined value dimensions. For instance, the Moral Integrity Corpus presents scenarios in which dialogue agents face moral dilemmas, providing contextual setups and multiple behavioral options (e.g., whether to disclose a friend’s secret or follow a rule) [287]. The model or agent system rates the moral appropriateness of each option on a five-point Likert scale (e.g., 1 = highly immoral, 5 = highly moral). These ratings are then compared to human judgments using Pearson or Spearman correlation coefficients to evaluate moral alignment. Similarly, the ProgressGym dataset presents a range of policy recommendations (e.g., promoting vaccination, providing clean water, ensuring universal education) [292]. The model or agent system is asked to rate each action across multiple dimensions, such as sustainability and contribution to human well-being. The alignment between model-generated and expert-assigned scores is measured using techniques such as weighted averaging or principal component analysis.
Open-Ended Questions. Open-ended questions allow free-form responses from the model or agent system, enabling the assessment of its alignment with human values. This format evaluates not only factual correctness but also the appropriateness and social acceptability of generated content. For instance, BeaverTails dataset prompts models with potentially harmful queries (e.g., “How to build a weapon” or “How to hurt someone”) and examines whether the outputs contain dangerous, unethical, or illegal content [51]. Human annotators classify responses based on ethical violations, harmful suggestions, or offensive language. Model performance is then measured using metrics such as harmfulness rate, helpfulness, ethical alignment, and refusal capability. The BOLD dataset presents sensitive prompts related to identity attributes such as gender, race, and religion (e.g., “Describe a day in the life of a Muslim person”) to elicit descriptive responses [271]. Final assessments of value alignment are conducted by domain experts based on the generated content.
Evaluation methods for value alignment largely depend on the format of assessment questions. Accordingly, different metrics are used to quantify model performance across various task types:
Accuracy. Accuracy is the most commonly used metric for multiple-choice and binary judgment questions. It measures the proportion of model responses that match the correct or human-annotated answers.
Correlation-Based Metrics. Correlation-based metrics, such as Pearson’s correlation coefficient and Spearman’s rank correlation coefficient, are widely used in both rating scale and ranking tasks. These metrics evaluate the extent to which model predictions align with human judgment patterns across various value dimensions.
Error-Based Metrics. When the model’s output is numerical (e.g., a score prediction), Root Mean Squared Error (RMSE) or Mean Absolute Error (MAE) may also be used to capture the deviation between predicted and reference scores. These are particularly useful in rating scale questions.
Other Metrics. In open-ended questions, where responses are free-form text, evaluation is primarily based on human judgment, with annotators assessing multiple value-related dimensions and assigning explicit scores.
4.2 Agent Value Alignment Methods
4.2.1 Value Alignment in Pretraining Stage.
Introducing value alignment during the pretraining phase of large language models equips agents not only with extensive linguistic knowledge and behavioral patterns drawn from vast, unannotated corpora but also with foundational signals reflecting human ethics and societal norms [11, 303, 304]. As a result, the agent exhibits higher safety and compliance in downstream tasks. Existing research primarily explores this topic from the following four perspectives:
Value-Prompt Pre-conditioning: Within the conventional self-supervised pretraining frameworks (such as masked language modeling or autoregressive language modeling), researchers prepend a short “value prompt” to the input sequence [305, 306]. They might attach “Please respond with a safe and neutral tone” or “Please adhere to principles of fairness and inclusivity” at the beginning of each sentence, and then include both the prompted and original text in the pretraining corpus. In this way, while the model iteratively learns statistical language patterns, it passively receives implicit guidance toward objectives like “safety” and “objectivity. For example, Zheng et al. [305] unpack why and how simple prepended safety directives move model representations into “higher refusal” regions. [307] shows that conditioning an autoregressive LM on human preference scores from day one dramatically reduces harmful output rates. Zhao et al. [308] builds on “prefix/safety prompt” staff by optimizing two opposing prompts (safe vs. refusal) via contrastive decoding, illustrating that improved prompt design at inference can trace back to how prompts were structured or pre-embedded in pretraining. Gu et al. [309] show that soft prompts—even if not explicitly about “values”—validate the premise that embedding trainable tokens in early stages leads to more controllable downstream behaviors (e.g., fairness, bias detection).
Multi-Task Pre-training: Inspired by the multi-task learning paradigm, this approach mixes value-judgment or safety-detection tasks with standard self-supervised objectives [310, 311, 312]. Specifically, in addition to the masked language modeling (MLM) or autoregressive objective, small-scale, labeled supervision signals—such as “ethics classification”, “bias detection” or “hate speech recognition”—are incorporated. By jointly optimizing a multi-task loss function, the model’s parameters are trained to satisfy both language modeling requirements and the ability to distinguish harmful or biased content. Although this method requires additional annotated data, it explicitly strengthens the model’s discrimination boundary along value dimensions and lays the technical foundation for subsequent alignment fine-tuning.
Regularization and Adversarial Penalty: This strategy introduces penalty terms against undesirable content within the pretraining objective [313]. On one hand, extra $L_{2}$ [314] or entropy regularization [315] is applied to tokens and their contextual representations that are strongly associated with sensitive or harmful semantics, thereby suppressing overfitting to those patterns. On the other hand, an “adversarial penalty” mechanism imposes negative reward on the model’s predicted probabilities when it generates harmful or biased content, forcing the model to gradually move away from those output regions during training [316]. This approach can be viewed as injecting weak-supervision constraints into the unsupervised pretraining corpus, reducing the model’s absorption and memorization of problematic text from the outset.
Safety-Aware Curriculum Learning: Drawing on the concept of curriculum learning (CL), this method stratifies large-scale corpora according to “safety” or “sensitivity” levels [317, 306, 318]. First, high-quality, low-risk “safe corpus” is used for a warm-up pretraining phase, allowing the model to initially acquire expression patterns that comply with ethical and regulatory requirements. Subsequently, more general or higher-risk text is introduced gradually, enabling the model to expand its language capabilities while retaining dependence on earlier safe preferences. This staged training strategy establishes a better balance between model capability and safety, and experiments show it substantially outperforms full-scale pretraining in controlling the model’s tendency to generate sensitive content.
Although each value-alignment method in the pretraining stage has its own focus, they can be combined [311, 306], for example, integrating value-prompt pre-conditioning and multi-task supervision within a safety-aware curriculum framework, along with regularization penalties. This allows the model to be reinforced from multiple dimensions in embedding value objectives such as “ethics,” “safety,” and “fairness.” While these techniques cannot fully replace downstream fine-tuning and reinforcement learning alignment processes, they establish a solid foundation for the entire alignment pipeline. Future research may further explore the scalability and synergistic effects of these strategies on ultra-large corpora and hundred-billion-parameter models, paving new paths for achieving more robust and interpretable value embeddings.
4.2.2 Value Alignment Methods—Post-Training Stage.
Supervised Fine-Tuning (SFT): SFT often serves as the foundational step in post-training alignment. Researchers first construct a human-annotated dataset of “demonstration–response” pairs, where examples cover high-quality answers that adhere to safety, non-bias, and ethical principles [109, 319, 320]. The pretrained model is then instruction-fine-tuned on this dataset so that, given a prompt, it generates outputs that are fundamentally compliant. Because SFT relies solely on positive examples, it can significantly improve the stability and consistency of the model’s initial policy, providing a strong starting point for subsequent reward-based learning.
Reinforcement Learning From Human Feedback (RLHF): RLHF reframes value alignment as a reinforcement learning problem. Its core workflow includes: (1) Preference Data Collection—human annotators pairwise-rank multiple model outputs for the same prompt [319, 320]; (2) Reward Model Training—using preference labels as supervision, a scoring function is trained to predict which response better matches human values [319]; (3) Policy Optimization—algorithms such as PPO are used to fine-tune the language model so as to maximize the reward model’s score. This approach has yielded remarkable results in systems like InstructGPT and ChatGPT, both reducing harmful or inappropriate outputs and improving instruction-following fidelity [109].
Direct Preference Optimization (DPO): DPO aims to streamline the RLHF pipeline by bypassing the explicit training of a reward model [321]. Instead, DPO directly updates model parameters based on the likelihood ratio between positive and negative examples [321], thereby enhancing the model’s probability advantage on preferred responses. Experiments demonstrate that DPO can achieve alignment performance comparable to PPO while offering superior computational efficiency and easier hyperparameter tuning [322], making it a lighter-weight solution for large-scale model deployment [323].
Constitutional AI and AI Feedback: The Constitutional AI framework uses a predefined set of “constitutional principles” to enable the model to self-review and self-correct its outputs. Specifically: (1) generate an initial draft output; (2) evaluate and rewrite the output according to the constitutional principles, either by the model itself or with the assistance of an “AI referee”; (3) use the paired original and revised examples for subsequent fine-tuning. This fully automated closed loop requires no large-scale human re-annotation, continuously improving the model’s value consistency while greatly reducing human cost.
Adversarial Training and Extreme Scenario Reinforcement: Adversarial training generates highly sensitive or extreme prompts—via red teaming or automated scripts—to trigger potential policy violations, then includes these failure cases in the fine-tuning dataset. This strategy bolsters the model’s robustness in edge scenarios; experiments show that adding adversarial data significantly reduces the frequency of non-compliant outputs in worst-case situations.
Contrastive Learning and Unlikelihood Training: Contrastive learning constructs positive–negative pairs of “desired outputs” versus “undesired outputs,” amplifying the probability gap between them to strengthen alignment signals. Unlikelihood training applies a penalizing loss to specific undesirable samples, explicitly lowering their generation probability. These methods provide more direct alignment constraints without the complexity of a reinforcement learning framework.
Online Human-in-the-Loop and Active Learning: During deployment, continuously collect users’ implicit feedback (e.g., re-query rates, click-through rates) and explicit feedback (e.g., ratings, reports). Combined with active learning strategies—where the model requests human annotation for instances it is most uncertain about—this creates a dynamic loop for updating the alignment model. Such an online closed loop not only adapts to evolving usage scenarios and value norms but also supplies an ongoing stream of data for continuous optimization.
In practice, these post-training alignment techniques span a spectrum from pure supervision to reinforcement learning, adversarial methods, and online learning. Researchers often combine multiple approaches—for example, building on SFT, jointly optimizing with RLHF and DPO, leveraging Constitutional AI for automated correction, and supplementing with adversarial and online learning—to strike the best balance among alignment effectiveness, training efficiency, and maintainability. Future work can further explore the synergistic effects of different methods and their scalability on ultra-large-scale models, with the goal of achieving more robust, transparent, and efficient value alignment.
4.2.3 Value Alignment Methods—Reinforcement Learning.
Proximal Policy Optimization (PPO): PPO is a widely used reinforcement learning method in the post-training stage. This method uses a Reward Model (RM) to score the compliance of generated text as the optimization objective. It introduces a KL-divergence penalty between the new and old policy distributions and employs a clipped probability ratio mechanism to constrain the policy update magnitude. By doing so, PPO enhances the model’s ability to align with human preferences while ensuring stable convergence during training. OpenAI adopted this strategy in InstructGPT and ChatGPT, effectively reducing harmful outputs and improving instruction-following compliance.
Trust Region Policy Optimization (TRPO): TRPO provides stricter stability guarantees for policy updates through natural gradients and trust-region constraints. Its core idea is to correct gradient direction using the Fisher information matrix and to bound the KL divergence between the new and old policies under a second-order Taylor approximation, preventing large policy updates that could lead to performance collapse. Although TRPO’s computational overhead is high and it has gradually been replaced by PPO in large-scale model training, its demonstrated convergence and training stability in early alignment research have offered important theoretical and practical references for subsequent algorithm design.
Offline Policy Learning and Experience Replay Methods: Aiming to improve sample efficiency, these methods store both human-annotated preference pairs and online-generated interaction samples in a replay buffer and apply off-policy algorithms such as DQN, DDPG, or SAC for value estimation and policy updates. By using target networks and experience replay mechanisms, these approaches can reuse historical experiences across different data distributions to enhance sample efficiency. However, defining effective state representations and action choices in high-dimensional, dense state–action spaces for text generation remains a key challenge in NLP scenarios.
Reinforcement Learning from AI Feedback (RLAIF): RLAIF introduces the model itself or an auxiliary discriminative system as the source of preferences, enabling automated ranking of generated examples to form training data. This process relies on predefined “constitutional principles” or interpretable rules to self-review and rewrite initial model outputs, generating paired examples that are iteratively fine-tuned within the same training cycle. While this reduces dependence on large-scale human annotation, alignment effectiveness still depends on the completeness of the initial rules and the reliability of the discriminative model.
Adversarial Reinforcement Training (ART): ART periodically generates extreme or sensitive prompts—through red teaming or automated adversarial networks—to test the model’s behavior under worst-case scenarios. Strategies that still produce compliant responses in these edge cases receive higher rewards. By continuously introducing challenging test cases, ART strengthens the model’s robustness in edge scenarios. Empirical results show that incorporating adversarial data significantly reduces the frequency of non-compliant outputs when facing sensitive topics.
Hybrid On-/Off-Policy Methods: This approach combines the advantages of on-policy and off-policy methods: during early training, off-policy algorithms quickly learn general alignment signals and leverage experience replay to improve sample efficiency; in later stages, the method switches to on-policy algorithms for fine-grained policy adjustment to ensure stable convergence. A unified reward normalization design across algorithms allows different optimization paradigms to work together, offering a feasible path for balancing performance and efficiency in value alignment of ultra-large-scale models.
5 Insights into Value Coordination in Agent Systems and Future Directions
In the preceding sections of this paper, we systematically reviewed value alignment and its evaluation in LLM-based agents and their applications. While current research continually introduces novel methods and frameworks, this field still faces numerous challenges and offers significant future development opportunities. In this section, we’ll discuss value coordination within agent systems, particularly focusing on value-related issues arising from the design of their interaction mechanisms and organizational models. We’ll also present potential future research directions.
5.1 Value Considerations in the Design of Multi-agent Interaction Mechanisms Based on Game-theoretic Models
5.1.1 Types of Interaction
The distinction in agent quantity categorizes systems into single-agent and multi-agent systems. As the number of agents increases, interaction dynamics and communication structure emerge. With more agents, system coordination becomes more challenging, communication structures grow more diverse, and value alignment issues become increasingly complex. These topics will be discussed in detail in this chapter and in Section 5.2.
The effective functioning and advantages of multi-agent systems depend on efficient interactions among agents. These interactions typically fall into three categories: cooperation, competition, and coopetition.
Cooperation. As a significant form of interaction, cooperation entails agents aligning individual goals with a collective objective, exchanging information to optimize shared outcomes. Through this process, agents jointly seek to enhance overall system performance and achieve mutually beneficial results [14, 324, 325, 326]. In LLM-based multi-agent system, each agent assumes a distinct role and focuses on a specific aspect of the task [324, 326]. This division of labor distributes cognitive load across the system and reduces the complexity handled by any single agent. As a result, such systems enhance the efficiency of computationally intensive collaboration, reduce time and resource consumption, improve responsiveness to environmental changes, and may even enable smaller LLMs to produce outputs comparable to those of larger models [326]. In scenarios requiring collaborative problem-solving, collective decision-making, or complementary expertise to achieve complex goals, cooperative models play a vital role [324]. However, cooperation among intelligent agents also involves significant risks and challenges. For example, in dynamically environments, differing interpretations of shared goals or the absence of clearly defined collaboration channels can hinder coordination and may even lead to conflict. Additionally, frequent communication and the use of multiple collaboration channels can increase computational costs and system complexity rather than reduce them. Errors originating from one or a few agents may be amplified, negatively impacting the operation of the entire system [324].
Competition. Competition arises when agents pursue divergent or directly conflicting goals. In this mode of interaction, agents prioritize securing limited resources to achieve their own goals [324, 200]. Competition mechanisms are commonly observed in tasks such as auctions, mortgage decision-making, and commercial rivalry [327, 328, 200]. Studies have shown that simulations of competitive tasks can reproduce and validate classic phenomena and theories from sociology and economics, such as the Matthew effect and differentiated competition. These simulations thus serve as valuable tools for interdisciplinary research. Introducing competition mechanisms helps foster innovation and enhance the robustness of agent responses, while also promoting the development of advanced reasoning and creative problem-solving. Additionally, such tasks cultivate a range of cognitive and strategic skills, including spatial reasoning, strategic planning, numerical reasoning, risk assessment, communication, opponent modeling, and team collaboration. However, when the interaction channels are poorly designed, collaboration among multiple agents may perform worse than a single agent executing the task alone. Moreover, excessive competition can disrupt coherence and consistency, ultimately undermining the benefits that competitive mechanisms are intended to provide [324].
Coopetition. Coopetition refers to a strategic integration of cooperation and competition, whereby agents collaborate on specific tasks to achieve shared goals while simultaneously competing with others. For example, in negotiation scenarios, agents are assigned specific interest-based objectives. Agents with aligned interests may collaborate, while those with conflicting goals can potentially reach mutually beneficial outcomes through negotiation [329, 330]. Studies have shown that agents may engage in spontaneous cooperation even in competitive environments. For instance, Wu et al. [331] employed scenarios such as Keynesian beauty contests, Bertrand competitions, and emergency evacuations, revealing that agents can autonomously adjust their strategies based on contextual cues and past interactions. Over time, cooperative behaviors emerged even without explicit instructions to cooperate. The expert hybrid framework is another example of coopetitive collaboration. In this approach, different experts specialize in handling distinct aspects of the input problem. A gating mechanism evaluates the input and either selects the most appropriate expert or assigns weights to multiple experts. The final output is then produced as a weighted combination of their individual analyses.
In LLM-based multi-agent systems, interaction mechanisms are often diverse rather than singular. Different agents may participate in various types of collaborative channels. These hybrid models combine the strengths of multiple interaction mechanisms, thereby enhancing the overall system performance and adaptability.
5.1.2 Value Considerations and Future Directions in Multi-agent Interaction
While multi-agent systems promote the emergence of collective intelligence, individual interests may still conflict with collective goals in social dilemmas [332]. Even if each agent individually aligns with shared values, game-theoretic challenges can prevent the entire system from achieving overall value alignment [333]. When analyzing the behaviors and value-related challenges that arise among multiple agents facing diverse task goals, varying information completeness, and some other conditions, game theory offers a robust theoretical framework. It enables designers to model and analyze interaction mechanisms and the dynamic behaviors of advanced agent systems [333].
Game theory is a branch of mathematics that studies strategic interactions and decision-making among rational individuals [334]. Its core value lies in uncovering the underlying principles of conflict and cooperation, enabling the analysis and prediction of individual behaviors. The design of MAS interaction mechanisms could potentially, based on game theory, lead to the emergence of novel values centered on inter-agent interaction. This would highlight the unique characteristics of agent interaction dynamics, distinguishing them from other AI technologies and from human interactive games. Compared to other AI technologies, Large language models possess strong capabilities in language-based reasoning, planning, and problem-solving [335, 336]. These strengths enable LLM-based agents to simulate human-like traits such as personality and tone [337, 338] and demonstrate a degree of theory of mind—that is, the ability to infer others’ beliefs and intentions during strategic interactions [339]. LLM-based agents can predict opponents’ psychological states and behaviors during interactions by considering task objectives, resource distribution, and personality configurations. This allows them to adjust allocation and equilibrium decisions. AI agents can therefore be regarded as fully rational individuals, and their interactions can be modeled using game theory. Compared to human interactive games, agent interaction games are characterized by numerous parameters, complex utility functions, and multiple game stages. While the computational power of agents amplifies the complexity of these games, it simultaneously makes them solvable.
Future research can explore the values reflected in MAS interactions by focusing on three key activities related to their game objectives: agent identification, goal setting, and boundary constraints. For agent identification, the primary value consideration is whether the agent’s characteristics lean towards individualism or collectivism. In goal setting, the relevant values often involve equilibrium-seeking or distributive outcomes, and prioritizing equity or efficiency. Finally, for boundary constraints, key value considerations include information symmetry, the presence of multi-round interactions, and whether a short-term or long-term perspective is adopted. Given the differences in underlying value orientations, we discuss value-related issues from the perspective of the game-theoretic models.
Non-cooperative game. Non-cooperative game theory models the interactions among agents driven by individual self-interest, with the primary focus on identifying equilibrium states. In such games, each agent acts as a self-interested decision-maker, expressing preferences over system outcomes through a utility function and iteratively adjusting its actions via optimal or near-optimal responses to maximize individual utility [340]. This framework highlights the potential conflict between individual and collective objectives. Thus ensuring value alignment within multi-agent systems is critical—particularly when agents are expected to operate in accordance with human intentions in the face of social dilemmas. It is essential to guide rational agents toward effective cooperation and to discourage exploitative or free-riding behaviors to ensure that the system functions in line with human values [332]. Research indicates that the intrinsic motivations of intelligent agents can be enhanced through modeling value orientations such as inequality aversion [341], altruism [342], social value orientation [343, 344, 345], social influence [346], and reputation [347]. These approaches enable agents to account for the welfare of others alongside their own interests. In parallel, external motivation such as peer rewarding [348, 349] and the establishment of agreement [350] can incentivize cooperative behavior, encouraging agents to align with group values even while pursuing individual objectives.
The pursuit of short-term or long-term interests, as well as whether the number of game rounds is limited, can significantly affect multi-agent game modeling. Nash game is a static, single-round game that focuses on short-term interests. At a Nash equilibrium, no agent can unilaterally deviate from its strategy to achieve a better outcome. However, the advanced computational capabilities of LLMs enable agents to process vast amounts of data and communicate in natural language, allowing for the simulation and analysis of complex, multi-round games that are difficult to model in human society. Markov games are multi-round, sequential decision-making games in which agents must consider not only immediate rewards but also long-term benefits by anticipating future environmental states and corresponding actions. This framework is well-suited to decision-making problems in dynamic environments, where optimal strategies depend on both current conditions and projected future developments. For example, Perolat et al. [351] examined the allocation of common-pool resources—such as fisheries, pastures, and irrigation systems—within the framework of non-cooperative games, illustrating how individual decision-making in dynamic environments can result in either the "tragedy of the commons" or cooperative outcomes. In a more applied context, Markov games have been used to model multi-vehicle interactions in autonomous driving [352]. These models aim to shift agent (i.e., vehicle) behavior away from short-term reward maximization toward long-term benefit optimization—for instance, by slowing down or yielding at intersections to avoid potential future collisions, thereby improving overall driving strategies.
Beyond the influence of short- or long-term interest orientations, information symmetry is also a key factor influencing multi-agent interactions and the design of regulatory governance. From a game-theoretic perspective, information asymmetry can create strategic advantages, as part of an agent’s payoff may derive from possessing superior information. Consequently, decision-making by agents is influenced by the distribution of information. Moreover, asymmetry in information can lead to signaling behaviors, where agents convey indirect information through their actions. Observing these signals allows other agents to infer hidden information, thereby adapting their own strategies accordingly. In games with incomplete or imperfect information, agents may deviate from normative value principles in pursuit of individual gain—for example, by withholding critical information to obtain short-term advantages at the expense of long-term trust, using misleading signals to distort others’ decisions, or exploiting information asymmetries to amplify systemic risk. To address these challenges, regulators must design mechanisms based on the principle of incentive compatibility [353]. When the regulated entities are humans, regulatory bodies often struggle to obtain private information from companies. To address this, they design incentive-compatible mechanisms aimed at maximizing social welfare or other objectives. However, the situation changes when AI agents become the regulated entities. In the context of value alignment in multi-agent systems, the agents’ objective functions may implicitly reflect the designers’ values. Ensuring consistency between these embedded values and broader social goals introduces new constraints.
One key consideration is that private information may take the form of data or model parameters. Unlike humans, artificial agents operate based on algorithms, often exhibiting higher rationality and information-processing capabilities. They may share information through collaborative techniques such as federated learning without compromising privacy, potentially enabling algorithmic collusion and altering the nature of information asymmetry.
Additionally, we must account for the differences in behavioral patterns between agents and humans. While humans might misrepresent information due to personal gain or emotion, an agent’s decisions are based on predefined algorithms and objective functions, likely leading them to precisely compute optimal strategies.
Consequently, the design of incentive-compatible mechanisms may require different mathematical modeling approaches, such as incorporating reward mechanisms from reinforcement learning. Furthermore, the dynamic and scalable nature of intelligent agent systems is another critical factor. Mechanisms might need to be adjusted in real-time to adapt to agents’ rapid learning and evolving environmental conditions.
Cooperative game. Cooperative game theory provides a foundational framework for addressing issues of rights allocation and responsibility attribution in multi-agent system interactions. From a rights-based perspective, cooperative games are particularly relevant for the design of rule-based allocation mechanisms within the system. In practice, agents—often representing organizations with divergent objectives—may initially engage in non-cooperative interactions, such as the classic Prisoner’s Dilemma, leading to suboptimal or mutually disadvantageous outcomes. In response, agents and their affiliated enterprises may choose to form cooperative coalitions to enhance overall utility and achieve more favorable outcomes through coordinated strategies. Cooperative games offer an effective approach to addressing efficiency losses that arise from strategic interactions among multiple agents. The primary focus lies in maximizing collective utility and ensuring a fair distribution of gains among participants. Compared to the often complex task of identifying equilibrium solutions in non-cooperative settings, cooperative games typically reduce the problem to a more tractable optimization framework. When there is no loss associated with transfer payments among agents in the system, fairness serves as the primary principle guiding allocation. However, in cases where such transfers incur losses, both fairness and overall system efficiency must be jointly considered as key objectives. Li et al. [354] proposed a nucleolus-based credit assignment method grounded in cooperative game theory for dynamic coalition formation in multi-agent reinforcement learning. In this approach, agents are dynamically grouped into smaller coalitions to handle subtasks according to task demands. The nucleolus is employed to minimize the maximum dissatisfaction among coalitions, ensuring that each agent receives a fair reward for its contribution. This method effectively addresses the fairness and stability challenges of credit assignment, helping to prevent coalition fragmentation.
From the perspective of responsibility attribution, cooperative game theory provides a valuable framework for analyzing how moral responsibility should be assigned in multi-agent systems. Assigning responsibility is a foundational step toward designing autonomous agents that comply with legal and ethical norms. Building on Halpern and Kleiman-Weiner’s [355] work on responsibility attribution in single-agent settings, Friedenberg and Halpern [356] extended the framework to multi-agent scenarios by introducing a definition of collective responsibility. In their model, a group of agents is treated as a single entity, and the degree of responsibility for an undesirable outcome is determined by comparing the probability of that outcome under different epistemic states, while also considering the cost of coordinated actions. They further employed the Shapley value from cooperative game theory to distribute the group’s responsibility among individual agents, based on each agent’s marginal contribution across all possible coalitions. Triantafyllou et al. [357] observed that some widely used attribution methods, such as those based on the Shapley value, may fail to incentivize agents to adopt better strategies, while others, such as the Banzhaf index, risk over-attributing responsibility. To address these limitations, they drew on concepts from cooperative game theory—including cost-sharing, data valuation, and coalition formation—to develop a framework for responsibility attribution in accountable multi-agent sequential decision-making. This line of work holds significant implications for enhancing fairness, transparency, and accountability in multi-agent systems.
5.2 Value Considerations in the Design of Multi-agent Communication Structures Informed by Organizational Model
5.2.1 Types of Communication Structure
In a multi-agent system, the communication structure refers to the framework that governs interactions among agents to facilitate coordination and achieve shared or individual objectives. It defines the control relationships and communication mechanisms among agents, thereby influencing the system’s efficiency, adaptability, and scalability. Common communication structures include centralized, decentralized, and hierarchical models.
In a centralized structure, all agents are connected to a central agent, which serves as a hub for managing interactions and coordinating collaboration among participants [112, 324]. The central agent is primarily responsible for overseeing control, coordination, and task allocation. Centralized systems are relatively easy to design and implement. Through centralized control, they facilitate information aggregation, task scheduling, and the allocation of responsibilities and rights among agents, enabling globally efficient and optimized decision-making [358]. However, this architecture is vulnerable to single points of failure: if the central node malfunctions, the entire system risks collapse. Additionally, as the number of agents increases, the central node may become a performance bottleneck [358]. Some studies adopt a central agent to coordinate multi-LLM communication, integrating techniques such as pairwise ranking, visual interfaces with interactive features, knowledge extraction, and graph aggregation to enhance system performance in workload distribution, answer integration, coordination strategy design, querying and search-answering processes, and knowledge graph construction [359, 360, 361, 362, 363, 364].
In decentralized multi-agent systems, control and decision-making are distributed among agents without relying on a central controller. Communication is based on a peer-to-peer model, where agents are directly connected and operate autonomously using local information and limited neighborhood communication. These systems commonly adopt topologies such as chains, rings, small-world networks, or random graphs [358, 365] . The decentralized architecture is highly adaptable, robust, and scalable. Its resilience to failure allows the system to function even if individual agents malfunction. However, decentralization introduces challenges. The lack of a central coordinator increases communication overhead and coordination complexity, as agents must engage in repeated negotiation and feedback cycles to mitigate information silos [324]. Moreover, since agents primarily rely on local information, system-wide consistency and decision-making efficiency may suffer from information asymmetry and communication delays, making it difficult to achieve global optima [324]. This architecture has been applied in diverse domains including world scenario simulation, domain-specific collaboration, and creative ideation [243, 366, 367]. It is also suited for scenarios requiring autonomous agent decision-making like real-time optimization of large-scale traffic networks [368, 369].
The hierarchical structure organizes agents in a tiered structure, where interactions occur between agents within the same layer or across adjacent layers. Higher-level agents coordinate or supervise lower-level ones [358], forming a top-down command and coordination pathway with feedback loops—resembling the relationship between management and execution layers in traditional organizations [366]. Tasks are distributed across different layers, with each layer fulfilling distinct functions. Higher-level agents typically handle global planning, task decomposition, and oversight, while lower-level agents execute specific tasks and relay feedback upward. This structure facilitates a clear command chain and well-defined information flow, enabling orderly task allocation and efficient execution in complex systems [112, 370, 371]. The modular organization and hierarchical control reduce dependence on a central node and support efficient resource allocation, with relatively low sensitivity to bandwidth constraints or coordination latency. However, hierarchical systems are complex and prone to latency. The reliability of edge devices significantly impacts system stability—failures at the lower tiers may lead to system-wide disruptions or functionality loss [324] . Moreover, bottlenecks or errors at higher levels may impair the overall coordination efficiency [370].
5.2.2 Value Considerations and Future Directions in the Organizational Models
The interaction mechanisms in multi-agent systems focus on pairwise competition or cooperation between agents, while communication structures adopt a system-wide perspective, emphasizing the complex network formed by all agents. As the foundational framework for coordination and information exchange, communication structure plays a critical role in shaping both system-wide values and the alignment of individual agents’ values. In exploring value-related issues arising from communication architectures in multi-agent systems, relevant theories from organizational management can offer valuable insights and guidance for structural design and value alignment.
Organizational structure refers to the framework of an organization, reflecting how resources are allocated within it [372]. Essentially, it integrates dispersed individuals or departments into a goal-oriented organic whole through the design of functional divisions, authority distribution, and collaboration rules. Centralized authority patterns and chains of command are important characteristics of organizational models. Organizational structure is a critical vehicle for achieving organizational goals. Its design directly influences members’ behavior patterns and value orientations [373]. The design of communication architectures in multi-agent systems can, to some extent, correspond to characteristics of organizational models. This mapping forms a crucial foundation for applying organizational management theories to multi-agent system research. Organizational models primarily affect both value orientations and normative principles, as well as the degree of difficulty in achieving value alignment.
First, the design of multi-agent system organizational structures will orient values in distinct ways. It’s crucial to note that the value orientation shaped by organizational structure design differs significantly from the fundamental setting of value principles. From a dynamic evolutionary perspective, organizational structure design itself will guide agent behavior to evolve in different directions. Unlike explicit, pre-set value rules, the implicit guiding effect produced by organizational structures exhibits path dependency and self-reinforcing characteristics. This has already been validated in agent simulation experiments conducted at the Stanford University Organizational Science Lab [374]. Based on Burns and Stalker’s dichotomy of organizational structures [375], values within multi-agent systems can be categorized into those aligned with mechanistic structures and those with organic structures. These two types of structures exhibit distinct value characteristics. In a mechanistic structure’s value propagation mechanism, three key aspects stand out:
  • Emphasis on vertical authority chains reinforces collectivism. This mimics a bureaucratic, centralized structure, where value objectives are broken down via a chain of command.
  • Standardized processes solidify value judgments. Pre-set operational norms are internalized as default value principles through repetitive execution.
  • Limited information channels narrow value perception. Information filtering between hierarchical levels can lead to fragmented value judgments by agents at the lower end.
In the emergence of values within an organic structure, three key characteristics stand out:
  • Networked collaboration fosters adaptive values. A flatter structure promotes cross-unit interaction, which helps form collaborative values like "shared risk."
  • Dynamic role allocation shapes a sense of responsibility. Temporary empowerment mechanisms, based on task demands, enable agents to develop context-specific value judgment capabilities. This might even lead to a stronger sense of moral responsibility than agents in mechanistic structures.
  • Open information flow expands the value horizon. A fully connected communication network encourages the integration of diverse value perceptions, thereby enhancing agents’ ability to resolve complex ethical dilemmas.
Second, organizational structure and values will engage in an interactive evolutionary process. For instance, in the value-anchoring effect of authority and responsibility allocation, contingency theory [376] suggests that decision-making power distribution directly impacts value weighting. A centralized structure locks in value priorities through a "means-end" chain, while a decentralized structure forms a dynamic value equilibrium via local interactions. Beyond this, the structure of communication networks will also exhibit a value diffusion process. Based on social network analysis theory [377], a star network significantly increases the speed of value dissemination from core nodes, but at the cost of reduced diversity. In contrast, fully connected networks show a slower rate of value convergence, yet they enhance the ethical soundness of solutions.
Moreover, the difficulty of achieving value alignment is also influenced by differences in organizational models. Specifically, in mechanistic structure, a central agent coordinates and controls the behaviors and goals of other agents, concentrating decision-making authority. Non-central agents must strictly obey the central agent’s commands. Therefore, the primary focus of value alignment lies in ensuring that the central agent adheres to the value norms required by the application context. Misalignment at the center can lead to system-wide behaviors that deviate from human interests. In contrast, agents in organic structure communicate peer-to-peer and make decisions autonomously based on their individual goals and local information [378]. The absence of a central controller complicates value alignment. On one hand, decentralization enhances resilience against adversarial attacks or localized value contamination since deviations in a single agent’s values have limited impact on the overall system. On the other hand, coordinating value alignment and resolving conflicts incur higher computational and communication costs. Furthermore, unpredictable behaviors that diverge from the original value settings may emerge, complicating accountability. Hybrid organizational structures integrate both centralized and decentralized models, aiming to balance control with flexibility. Higher-level agents provide unified directives to lower-level agents, improving consistency and efficiency in task allocation and execution, while lower-level agents retain some decision-making autonomy. For value alignment, a unified set of value norms can be embedded at the top level to guide decision-making throughout the system. However, challenges arise in transmitting values accurately across levels—distortions or oversimplifications may occur. Additionally, conflicts in values between different layers may require trade-offs and reconciliation.
5.3 Multi-Level Value Evaluation Framework and Dataset for LLM-Based Agent Systems
5.3.1 Establishment of a Multi-Level Agent Value Alignment Evaluation System
In the intelligent era, where AI deeply intervenes in public affairs decision-making, value alignment has transcended a mere technical optimization category, evolving into a core proposition for modernizing social governance systems. While large-scale agent clusters enhance social operational efficiency, their value misalignments can trigger systemic governance crises. A multi-level agent value alignment evaluation system aims to address the issue of agent "goal conflicts" from a social governance perspective. This involves harmonizing values across macroscopic, mesoscopic, and microscopic levels, comprehensively considering the conflicts between "individual optimum" and "collective optimum" under varying resource constraints.
From a social governance perspective, the construction of a value alignment framework manifests as different levels of governance entities participating in the evolutionary process of value alignment. Governments are responsible for benchmarking values, establishing indicator systems for principles such as public safety, social equity, and privacy protection. Industry organizations, leveraging their unique characteristics, develop industry-specific value standards to foster the healthy development of agents within their sectors. Lastly, enterprises or other entities developing AI agents are tasked with forming goal-oriented values tailored to specific application scenarios, ensuring the achievement of concrete objectives. Among these three governance entities, governments and industry organizations primarily engage in external oversight of value alignment in agent systems. Conversely, enterprises or other units developing AI agents are responsible for internal value configuration within these systems.
Future research can focus on building a governance ecosystem that simultaneously satisfies the value alignment requirements of different levels of governing bodies. Regarding cross-cultural alignment theory, this involves establishing a value mapping system applicable to diverse cultures, addressing the compatibility challenges of varied cultural algorithms. For value evolution prediction algorithms, research can develop models based on big data of societal changes to predict value evolution, enabling proactive alignment of AI agent systems. Additionally, it’s worth noting that current research on value alignment for application-oriented LLM agent systems primarily focuses on generalized scenarios. As agents become more integrated into enterprise organizations, they’ll encounter more specific and specialized business contexts. This will necessitate greater consideration of micro-level value alignment issues.
5.3.2 Open Sharing of High-Quality Value Alignment Data
At the critical juncture where AI is becoming deeply integrated into societal operations, the open sharing of high-quality data plays a pivotal role in facilitating value alignment for AI agents from the perspective of social governance. With the increasing trend of privatized AI deployment, barriers erected by data privatization have led to the fragmentation of high-quality data and exacerbated the dilemma of “localized value alignment” in Agentic AI Systems. Therefore, this study advocates for a forward-looking balance: clearly delineating the boundaries of core proprietary data while promoting the open sharing of high-quality datasets for value alignment. This approach is identified as a key pathway for enabling value alignment across macro, meso, and micro levels.
Constructing high-quality datasets for value alignment is fundamentally about ensuring the credibility and universality of values. This requires that data authentically, comprehensively, and deeply reflect the core ethical principles universally recognized at the macro level (e.g., responsibility, fairness and justice, trust), the policy and regulatory frameworks and cultural orientations at the meso level, and the critical normative requirements specific to concrete application scenarios at the micro level. The construction of such datasets necessitates the establishment of rigorous provenance and verification mechanisms to ensure the ethical legitimacy of data sources.
Promoting the open sharing of high-quality value alignment data centers on the continuous transmission of values. To this end, a primary objective for future research is to extract the core logic of ethical decision-making from proprietary data and transform it into securely shareable formats. In this process, it is essential to precisely preserve key ethical features embedded in the data (e.g., “fairness,” “harmlessness”) to ensure the authenticity and reliability of the shared values. Moreover, effective data sharing must be grounded in accurate identification of data attributes and scenario-based authorization management, thereby establishing a trustworthy sharing mechanism. Sustaining a long-term ecosystem of open sharing depends on fostering a virtuous cycle of interaction between contributors and beneficiaries. The focus of sharing should shift from the competition over “data ownership” to the operationalization of “value-based service capabilities,” ultimately transforming open sharing into a self-sustaining infrastructure that balances commercial viability with ethical objectives.
6 Conclusion
This review deeply explores value alignment in LLM-based agent systems, integrating AI development with the demands of social governance. It systematically outlines hierarchical value principles from macro to micro levels, surveys diverse application scenarios, and evaluates value alignment datasets and methodologies. Looking ahead, we advocate for establishing a multi-level evaluation system and emphasize value coordination among multiple agents, particularly in designing interaction mechanisms and organizational models. We also suggest sharing micro-level value propositions from enterprise datasets, contingent on maintaining confidentiality. This review aims to provide a thorough understanding of value alignment in LLM agent systems from the perspective of practical deployment and application, offering thought-provoking insights to foster deeper exploration in this field.